### v1

# Complements of discriminants of real parabolic function singularities. II

Abstract.
We provide the complete list of the connected components of the spaces of non-discriminant functions within standard versal deformations of function singularities of classes $X_{9}$ and $J_{10}$. Thus, we prove (and improve in one particular case) all corresponding conjectures from the previous work [12] with the same title.
Key words and phrases:
discriminant, singularity, surgery, versal deformation, isotopy classification
1991 Mathematics Subject Classification:
Primary: 14Q30. Secondary: 14B07, 14P25
1. Introduction
Discriminant varieties of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of these functions are singular. These varieties appear in many problems of PDE theory and integral geometry (often under the name of wavefronts), see e.g. [1], [2], [3], [10], [11].
Many important functions in these theories are defined by integral transforms and are regular outside the discriminant varieties of deformations of corresponding generating functions. Their analytic behavior near the singular points of the discriminants depends on the singularity type of these functions and on the connected component of the complement of the discriminant. Thus, enumerating these components is one of the first steps in the study of their analytical properties.
The list of function singularities begins with simple singularities, which were distinguished by V. Arnold, see [2]. E. Looijenga [8] implicitly listed the local connected components of complements of their discriminants. O. Viro [18] listed all possible topological types of non-discriminant functions representing these components. (By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood. The topological type is an invariant of the components of the discriminant complements.) In [13], it was proven that each topological type of this list represents exactly one connected component of the discriminant complement.
The next important class of function singularities are the parabolic ones, which were also distinguished by V. Arnold. The classification of complex parabolic singularities of functions consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$; the classification of their real forms has more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$.
In [12], a system of invariants of the connected components of the complements of the discriminant varieties of real function singularities was described. The complete lists of the possible values of these invariants for parabolic singularities were found, and all of these values were realized by concrete polynomials. We also conjectured that the resulting lists of connected components thus found in [12] are complete. These lists consist of 7, 14, 52, 11, and 33 components for the singularities of classes $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two mutually symmetric sets of perturbations of the $J_{10}^{1}$ singularities indicated in [12] consist of pairs of different connected components. Thus, the number 11 above is replaced by 13. These results are summarized in Theorems 5, 13, 18, 28, and 31 below. The classes $P_{8}^{*}$ are not considered yet.
After the publication of [12] and [13], I discovered that O. Viro [18] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities (as well as of simple singularities).
Table 1. Numbers of isotopy and topological classes of non-discriminant perturbations of $X_{9}$ and $J_{10}$ singularities
Singularity class $X_{9}^{\pm}$ $X_{9}^{1}$ $X_{9}^{2}$ $J_{10}^{1}$ $J_{10}^{3}$
Topology 6 12 50 10 31
Isotopy 7 14 52 13 33
It turns out that the “isotopy” classification of parabolic singularity perturbations considered in the present work is only slightly stricter than the topological classification, see Table 1.
E. Looijenga [8] proved that all local components of discriminant complements of simple singularities are contractible. Conversely, we show that at least some three such components of parabolic singularities of classes $X_{9}^{+}$ and $X_{9}^{-}$ have infinite one-dimensional homology groups, see Theorem 12 on page 12.
An application to hyperbolic PDE’s. In the theory of hyperbolic PDE’s, the local Petrovskii lacunas are the local connected components of the complements of wavefronts, in which the wave functions behave regularly, see e.g. [1], [10].
In [11], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that this list is complete. The enumeration of all local components of wavefront complements given below proves this conjecture for all parabolic singularities of classes $X_{9}$ and $J_{10}$, see § 4.
2. Results for $X_{9}^{*}$ classes
Table 2. Normal forms of $X_{9}$ and $J_{10}$ singularities in two variables
Notation Normal form Restriction
$X_{9}^{+}$ $x^{4}+2Ax^{2}y^{2}+y^{4}$ $|A|<1$
$X_{9}^{1}$ $xy(x^{2}+2Axy+y^{2})$ $|A|<1$
$X_{9}^{2}$ $x^{4}+2Ax^{2}y^{2}+y^{4}$ A<1A<-1
$J_{10}^{1}$ $(x^{2}+y^{2})(x-\gamma y^{2})$ $\gamma\in(-\infty,\infty)$
$J_{10}^{3}$ $(x^{2}-y^{2})(x-\gamma y^{2})$ $\gamma\in(-1,1)$
Definition 1.
A function singularity is a germ of a $C^{\infty}$-smooth function ${\mathbb{R}}^{2}\to{\mathbb{R}}$ at its critical point.
Real parabolic function singularities of type $X_{9}$ form four families $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$. Any function singularity of classes $X_{9}^{+}$, $X_{9}^{1}$ or $X_{9}^{2}$ has the normal form indicated in Table 2 in appropriate local coordinates centered at the critical point. The class $X_{9}^{-}$ (and all considerations concerning it) can be reduced to the case $X_{9}^{+}$ by multiplying by 1-1.
The standard miniversal deformations of functions of all three classes $X_{9}^{+}$, $X_{9}^{1}$, and $X_{9}^{2}$ have the same form: they are families of polynomials
(1) $f_{A}(x)+\lambda_{1}+\lambda_{2}x+\lambda_{3}y+\lambda_{4}x^{2}+\lambda_{5}xy+\lambda_{6}y^{2}+\lambda_{7}x^{2}y+\lambda_{8}xy^{2},$
depending on nine parameters $\lambda_{1},\dots,\lambda_{8},$ and $A$, where $f_{A}$ is the corresponding normal form from Table 2. By the versality theory (see, e.g., [2]), the study of all other deformations of these singularities can be reduced to these.
Definition 2.
A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (1) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ consists of four different lines in ${\mathbb{C}}^{2}$.
The spaces $\Phi_{+}$, $\Phi_{1}$, and $\Phi_{2}$ consist of all polynomials of the form (1) with principal parts of classes $X_{9}^{+}$, $X_{9}^{1}$, and $X_{9}^{2}$, respectively.
The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$. The discriminant subvarieties of these spaces consist of functions that have at least one real critical point with zero critical value. These varieties divide the spaces $\Phi_{*}$ into connected components. The enumeration of these components is our goal in this section.
In §2 of [12] we introduced the set-valued invariant of these components. Let us briefly review this construction.
The virtual function associated with a generic real polynomial is a collection of certain topological characteristics of the complex zero-level set of this polynomial, depending also on the collections of paths in ${\mathbb{C}}^{1}$ defining the bases of vanishing cycles in the homology of this level set. Namely, these data of a function $f(x,y)$ consist of
  • the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented basic vanishing cycles in the submanifold of ${\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
  • the string of intersection indices of these cycles with the set of real points,
  • the Morse indices of real critical points of $f$, and
  • the numbers of positive and negative critical values of $f$.
The (abstract) virtual functions of class $\Phi_{*}$ include all virtual functions associated with generic real polynomials of this class, as well as all analogous collections of data that can be obtained from them via standard flips that model surgeries of real functions. The formal graph of a singularity class $X_{9}^{*}$ is the graph, whose vertices correspond to all the virtual functions of the corresponding class $\Phi_{*}$, and whose edges correspond to the standard flips. The virtual components of the formal graph are the connected subgraphs, into which the formal graph splits when all edges modeling the surgeries crossing the discriminant are removed. The set-valued invariant of a generic polynomial $f$ of class $\Phi_{*}$ is the virtual component $S(f)$ of the formal graph containing the virtual functions associated with this polynomial.
These objects have the following obvious property.
Proposition 1.
If two polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$, S(f~)S($\tilde${f}) are the same. $\Box$
Enumerating the virtual components of a finite formal graph is a combinatorial problem that can be solved by a computer program. The work [12] is essentially a list of the results of these computations for all parabolic singularities and of their immediate consequences in the terms of real functions. The remaining problem is to find, for each virtual component, all real connected components with this value of the set-valued invariant.
It follows from the results of [5] that, for parabolic singularities, all standard flips of virtual functions can be realized by real surgeries of generic real functions. In particular, the following statement holds.
Proposition 2 (see [15], Proposition 3).
For any class of polynomials $\Phi_{+}$, $\Phi_{1}$, or $\Phi_{2}$,
a)) every virtual function of this class is associated with a generic real polynomial $\tilde{f}$ of the same class;
b)) every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial from the same connected component as $f$ of the complement of the discriminant. $\Box$
Remark 3.
In the statement of Proposition 3 of [15], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four polynomials with non-degenerate principal homogeneous parts. However, the entire construction proving this proposition can be performed within the spaces of the versal deformations (1). Thus, it also proves Proposition 2.
These considerations enable us to estimate the number of isotopy classes of generic Morse perturbations associated with any given virtual function. These estimates are given in Propositions 8, 15, and 20 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately. Consequently, we obtain estimates of the number of connected components of the complement of a real discriminant which are associated with each virtual component. Specifically, in the cases $X_{9}^{+}$ and $X_{9}^{2}$, all real connected components of discriminant complements associated with the same virtual component can be obtained from each other by the action of a transformation group of order eight. In the case $X_{9}^{1}$ the analogous group is of order four. Then, for each virtual component in the list from [12], we compare all real components represented by the polynomials associated with some its virtual function and forming an orbit of this group, and determine which of these components are identical. This comparison provides the results of the present work concerning the $X_{9}$ singularities. The strategy for the $J_{10}$ singularities is exactly the same.
Remark 4.
In [15], [16], and [17], we use the same strategy with a different definition of virtual components. In these works, we classify the Morse perturbations of function singularities up to isotopy within the space of Morse functions instead of the space of non-discriminant functions. Thus, the virtual components of the formal graph are determined there by removing the edges that change the Morse type.
Definition 3.
For any generic polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$, $\mbox{Ind}(f)$ is the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
It is easy to see that this number is an invariant of the connected components of the discriminant complements.
2.1. Computations for the $X_{9}^{+}$ singularities
In this subsection, we prove the following statement, conjectured in [12].
++$(a)$
(b)(b)++
(c)(c)++
++$(d)$
(e)(e)++
++$(f)$
++$(g)$
Figure 1. Perturbations of $X_{9}$ singularity
Theorem 5.
There are exactly seven connected components in the space of non-discriminant functions of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials representing these components are shown in Fig. 1.
Remark 6.
Pictures $(f)$ and $(g)$ of Fig. 1 are topologically equivalent. However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$. Figure 1 coincides with the well-known list of affine isotopy classes of bounded degree four algebraic plane curves. In our case, however, the isotopies are allowed only in a much smaller space $\Phi_{+}$. In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
The starting point of the proof of Theorem 5 is the following proposition.
Proposition 7 (see [12]).
There are exactly seven virtual components of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials associated with virtual functions representing these virtual components are shown in Fig. 1. $\Box$
Definition 4.
Denote by $G_{0}$ the group of linear transformations of ${\mathbb{R}}^{2}$ generated by reflections $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(y,x)$
Clearly, this group is isomorphic to the group of symmetries of the square, in particular is of order eight. This group acts on the space $\Phi_{+}$ of polynomials (1) with principal parts of class $X_{9}^{+}$ (each transformation $\Gamma:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ moves the polynomial $f$ to $f\circ\Gamma$). This action preserves the set of generic polynomials.
Proposition 8.
If two generic polynomials, $f$ and $\tilde{f}$, of class $\Phi_{+}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of class $\Phi_{+}$ containing $f$ and $\tilde{f}$ are mapped to each other by an element of the group $G_{0}$.
Definition 5.
For any polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree four with principal homogeneous part $x^{4}+2Ax^{2}y^{2}+y^{4}$, denote by ${\bf T}(f)$ the unique polynomial of the form (1) with the same principal part, which is obtained from $f$ by a parallel translation of ${\mathbb{R}}^{2}$. (This is the translation by the vector $\frac{1}{4}(a,b),$ where $a$ and $b$ are the coefficients of $f$ at the monomials $x^{3}$ and $y^{3}$).
Lemma 9 (see [15], Lemma 4).
A function of the form ((1)) with a positive principal homogeneous part has a critical point of class $E_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{27}{256}$ if and only if it is one of the four polynomials
(2) $\hskip 82.51282pt{\bf T}(x^{4}\pm x^{3}+y^{4})\qquad\mbox{ or }\qquad{\bf T}(x^{4}+y^{4}\pm y^{3})\ .\hskip 82.51282pt\Box$
Proof of Proposition 8. This proof almost coincides with the proof of Proposition 3 in §3.5 of [15]. In the next paragraph, we present the outline of this proof; we refer to [15] for all common details.
Using the Lyashko–Looijenga covering (in the form described in [5]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the discriminant and the set of not strictly Morse functions). Since the virtual function of $\tilde{f}$ is the same as of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that commutes with the Lyashko–Looijenga map: the map induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$. Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are the same, or they are mapped into each other by this reflection. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this reflection, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads to the previous situation. $\Box$
By Proposition 8, any virtual component mentioned in Proposition 7 is associated with at most eight real connected components. Let us prove that in fact each virtual component represented by a polynomial shown in Fig. 1 is associated with a single real connected component.
Pictures $(a)$, $(b)$, $(c)$, $(f)$ of Fig. 1 are realized, respectively, by four polynomials
$x^{4}+y^{4}+t,$
$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\hskip 23.49976pt\alpha\in(0,1/4),$
$x^{4}+y^{4}-t,$
$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(1/4,1/2),$
with arbitrary $t>0$. All of these polynomials are invariant under the entire group $G_{0}$.
Picture (d) can be represented by the polynomial
(3) $x^{4}+2Ax^{2}y^{2}+y^{4}+2x^{2}-2y^{2}+1/2$
with arbitrary $A\in(-1,1)$. This polynomial is invariant under the transformations $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y)$. Thus, its $G_{0}$-orbit consists only of itself and the polynomial
(4) $x^{4}+2Ax^{2}y^{2}+y^{4}-2x^{2}+2y^{2}+1/2$
The polynomials (3) and (4) are connected by the family of polynomials
(5) $x^{4}+2Ax^{2}y^{2}+y^{4}+2(x\cos\varphi+y\sin\varphi)^{2}-2(x\sin\varphi-y\cos\varphi)^{2}+1/2,$
depending on the parameter $\varphi\in[0,\pi/2]$.
Lemma 10.
If $A<1$ is sufficiently close to $1$, then all polynomials of this family are non-discriminant, in particular their zero-level sets are all smooth and isotopic to each other.
Proof. If we replace $A$ by 1 in (5), then all these polynomials will be obtained from each other by the rotations of ${\mathbb{R}}^{2}$ by angles $\varphi\in[0,\pi/2]$. Each of these polynomials has only three real Morse critical points: one at the origin and two on the unit circle. If $A=1-\varepsilon$ with sufficiently small $\varepsilon$, then the real zero-level set deforms only slightly in the domain $x^{2}+y^{2}\leq 2$ and maintains its topological structure in this domain for any $\varphi$. It also remains empty outside this domain. $\Box$
Thus, the connected component of the space of non-discriminant polynomials containing (3) is also invariant under the transformation $(x,y)\mapsto(y,x)$.
The topological type of the curve shown in picture $(e)$ can be realized by any of the four polynomials
(6) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(x^{3}-3xy^{2})+\varkappa)$
and
(7) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(y^{3}-3x^{2}y)+\varkappa)$
with arbitrary $A\in(-1,1)$ and sufficiently small $\varkappa>0$ (where the restriction on $\varkappa$ depends on $A$). These four polynomials form an orbit of the group $G_{0}$. All of them belong to the one-parameter family of polynomials
(8) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}+\mbox{Re}(e^{i\tau}(x+iy)^{3}))+\varkappa),$
$\tau\in[0,2\pi]$. For $A=1$, all these polynomials are obtained from each other by the rotations of ${\mathbb{R}}^{2}$. They are bounded from below, have a critical point of class $D_{4}^{-}$ with zero critical value at the origin and three real minimum points with critical value $\frac{27}{256}+\varkappa$, and no other critical points. Therefore, for $\varkappa<\frac{27}{256}$ all of them realize the topological picture $(e)$. Analogously to Lemma 10, taking $A<1$ very close to 1, we will not change these topological pictures for any $\tau$. So, the connected component of the discriminant complement containing these polynomials is invariant under the action of the group $G_{0}$.
The picture $(g)$ can be realized by either of the four polynomials
(9) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10x(x^{2}-3y^{2})+15(x^{2}+y^{2})-5)$
and
(10) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10y(y^{2}-3x^{2})+15(x^{2}+y^{2})-5).$
with sufficiently small $\varepsilon>0$. These four polynomials also form an orbit of the group $G_{0}$.
The family of polynomials
(11) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}+10\mbox{Re}\left(e^{i\tau}(x+iy))^{3}\right)+15(x^{2}+y^{2})-5),$
$\tau\in[0,2\pi],$ contains all four polynomials (9), (10) for $\tau=0,$ $\frac{\pi}{2}$, $\pi$, and $\frac{3\pi}{2}$. All elements of this family lie in the same connected component of the complement of the discriminant. This completes the proof of Theorem 5. $\Box$
Remark 11.
To realize the pictures ($d$), ($e$), and ($g$) by arbitrarily small perturbations of the $X_{9}^{+}$ singularities, we can multiply by $t^{4-k}$ all monomials of degree $k$ in the formulas (3), (6), and (9).
The 1-cohomology of discriminant complements. Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, that is, the space of $k$-point subsets of the plane. For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$. Let $C$ be a connected component of the discriminant complement in the space $\Phi_{+}$ consisting of polynomials with $k\geq 2$ non-nested ovals. Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(C)$ is well-defined as follows. For any such polynomial $f$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals. These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $C$. The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $C\to B({\mathbb{R}}^{2},k)$. Clearly, these maps defined by different cross-sections are homotopic and define the same homomorphism of cohomology rings.
Define the class $W\in H^{1}(C)$ as the class obtained by this map from the generator of the group $H^{1}(B({\mathbb{R}}^{2},k))$.
Theorem 12 (cf. [14], § 1.8.3).
The class $W$ takes all integer values on some 1-cycles in each of the connected components shown in pictures ((d)), ((e)), and ((g)) of Fig. 1.
Proof. These cycles are the families (5), (8), and (11). $\Box$
2.2. The components for the $X_{9}^{1}$ singularities
Define the group $G_{1}$ as the group of linear transformations of ${\mathbb{R}}^{2}$ generated by the symmetries $(x,y)\mapsto(y,x)$ and $(x,y)\mapsto(-x,-y)$. Clearly, this group is isomorphic to ${\mathbb{Z}}_{2}^{2}$. It acts on the space of polynomials $\Phi_{1}$ and preserves the set of generic polynomials. Polynomials obtained from each other by this action have equal associated virtual functions and so are not separated by the set-valued invariant.
Additionally, the involution
(12) $f(x,y)\mapsto-f(x,-y)$
acts on the space of polynomials (1) of class $\Phi_{1}$. It preserves the set of discriminant polynomials and hence permutes the connected components of its complement. This permutation has no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components related by this involution is always 1-1. This involution naturally extends to the set of associated virtual components: each virtual component associated with the polynomials from the connected component containing the polynomial $f(x,y)$ goes to a virtual component associated with the polynomials isotopic to f(x,y)-f(x,-y). The resulting involution on the set of virtual components of class $\Phi_{1}$ also has no fixed elements.
For any polynomial $f$ with a principal homogeneous part of the form
(13) $xy(x^{2}+2Axy+y^{2}),\qquad|A|<1,$
its zero-level set goes to the infinity along four branches: two vertical and two horizontal. If the polynomial is non-discriminant, then each vertical branch is matched with the horizontal branch belonging to the same component of the zero-level set. The involution (12) acts on the zero-level sets as the reflection in the line $\{y=0\}$. In particular, it always alters the matching of the vertical and horizontal branches.
++--++$(a)$
++--++$(b)$
++--++$(c)$
++--++$(d)$
++--++$(e)$
++--++$(f)$
++--++$(g)$
Figure 2. Perturbations of $X_{9}^{1}$ singularity
Theorem 13.
There are exactly 14 connected components of the complement of the discriminant in the space of polynomials ((1)) of class $\Phi_{1}$. The shapes of the zero-level sets of polynomials representing some seven of them are shown in Fig. 2, and the remaining seven can be obtained from these by the involution ((12)).
The proof of this theorem is based on the following two propositions.
Proposition 14 (see [12]).
There are exactly ten virtual components of type $X_{9}^{1}$. They split into five pairs of virtual components related to each other by the involution induced by ((12)). The topological types of zero-level sets of polynomials associated with elements of some five of the virtual components ((one from each pair)) are shown in pictures ((a)), ((b)), ((c)), ((d), and ((f)) of Fig. 2. $\Box$
Proposition 15.
If two generic polynomials $f$ and $\tilde{f}$ of class $\Phi_{1}$ are associated with the same virtual function, then their connected components in the space of generic polynomials are moved to each other by an element of the group $G_{1}$.
Definition 6.
For any polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree four with principal homogeneous part (13), denote by ${\bf\tilde{T}}(f)$ the unique polynomial of the form (1) with the same principal part, which is obtained from $f$ by a parallel translation of the plane. (This is the translation by the vector $(\alpha,\beta),$ where $\alpha$ and $\beta$ are the coefficients of $f$ at the monomials $y^{3}$ and $x^{3}$).
Lemma 16.
A polynomial of the form ((1)) with the principal part ((13)) has a critical point of class $E_{7}$ with zero critical value and a critical point of class $A_{2}$ with critical value $3^{\frac{15}{2}}2^{-8}$ if and only if it is one of the following four polynomials:
(14) ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm x^{3})$
or
(15) ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm y^{3})\ .$
Proof. Let $f$ be a polynomial that satisfies the conditions of this lemma. By Lemma 5 of [15], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (13). In the affine coordinates $(x,y)$ centered at the $E_{7}$ point, the monomials of degree $\leq 2$ of this polynomial vanish, and the homogeneous part of degree three has the form $ax^{3}$ or $ay^{3}$, $a\neq 0$. The critical value of this function at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$. This value is equal to $3^{\frac{15}{2}}2^{-8}$ if and only if this sign is ++ and $\alpha$ equals $1$ or 1-1. $\Box$
Proof of Proposition 15. The deduction of this proof from Lemma 16 essentially repeats the proof of Proposition 8 with one simplification. In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial involutions that commute with the Lyashko–Looijenga covering (see [15], § 4.5). Therefore, the statement follows immediately from the fact that all points (14) and (15) belong to the same orbit of the group $G_{1}$. $\Box$
Proof of Theorem 13. Proposition 15 implies that any virtual component of the formal graph of the $X_{9}^{1}$ singularity is associated with at most four real connected components mapped to each other by the group $G_{1}$. Below we prove that six of the virtual components described in Proposition 14 are associated with single real connected component each, while the four remaining virtual components are each associated with a pair of real connected components that are mapped into each other by the central symmetry in ${\mathbb{R}}^{2}$.
Pictures $(a)$, $(b)$, and $(c)$ of Fig. 2 are realized by the polynomials
$xy(x^{2}+y^{2})-t,$
$xy(x^{2}+y^{2})-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(0,1/2),$
and
$xy(x^{2}+y^{2}-t)-\alpha t^{2},\qquad\alpha\in(0,1/8)$
for any $t>0$. All of these polynomials are invariant under the action of the group $G_{1}$. The polynomials obtained from them by the involution (12) and representing other three connected components of the discriminant complement are also invariant under this action.
Picture $(d)$ (respectively, $(e)$) is realized by the polynomial
(16) ${\bf\tilde{T}}\left((x^{2}+y^{2}-t^{2})(x+0.9t)(y+0.9t)+\alpha t^{4}\right),$
$\alpha\in(0,0.003)$, with positive (respectively, negative) $t$. These two pictures are topologically equivalent. However, it easily follows from Bezout’s theorem that these polynomials represent different connected components of the discriminant complement, see [15].
Each polynomial (16) is invariant under the transformation $(x,y)\mapsto(y,x)$. The transformation $(x,y)\mapsto(-x,-y)$ maps each such polynomial to the polynomial with the opposite value of parameter $t$. In particular, it permutes the connected components described by pictures $(d)$ and $(e)$.
Pictures $(f)$ and $(g)$ of Fig. 2 are realized by polynomials
(17) $xy\left((x-2t)^{2}+(y-2t)^{2}-t^{2}\right)-t^{4}=0$
with positive and negative values of $t$, respectively. Each of these polynomials is invariant under the transformation $(x,y)\mapsto(y,x)$; the transformation $(x,y)\mapsto(-x,-y)$ maps each polynomial (17) to the polynomial with the opposite value of the parameter $t$, thus it permutes two connected components represented by these polynomials.
The same considerations hold for components obtained from these by the involution (12). This concludes the proof of Theorem 13. $\Box$
2.3. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.1) acts on the space $\Phi_{2}$ of all polynomials (1) with A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way. All components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
In addition, the operation
(18) $f(x,y)\mapsto-f\left(\frac{x+y}{\sqrt{2}},\frac{x-y}{\sqrt{2}}\right)$
acts on the space $\Phi_{2}$ (and on the entire space of polynomials of degree four). This operation always alters the real connected components of the discriminant complement, as well as the associated virtual components of type $\Phi_{2}$. Indeed, it maps any generic polynomial with value $I$ of the invariant Ind to a polynomial with value 3I-3-I. This operation is of order eight. All even iterations of it act as rotations by angles of $\frac{\pi k}{2}$ and thus belong to the group $G_{0}$. So, the entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation. The cardinalities of all these orbits are divisors of 8.
----++$(a)$
----++$(b)$
----++$(c)$
----++++$(d)$
----++++$(e)$
----++++$(f)$
----++++$(g)$
----++++$(h)$
++++++--$(i)$
Figure 3. Non-discriminant perturbations of $X_{9}^{2}$
Proposition 17 (see [12]).
There are exactly 18 virtual components of class $\Phi_{2}$. The shapes of the zero-level sets of the polynomials representing nine of these virtual components are shown in Fig. 3 (see also Fig. 33 in [18]), and the polynomials representing the remaining nine virtual components are obtained from these by the operation ((18)).
Theorem 18.
There are exactly 52 connected components of the discriminant complement in the space $\Phi_{2}$.
Specifically, the virtual components represented by pictures $(a)$ and $(b)$ of Fig. 3 are associated with single real connected components of the discriminant complement; the virtual components represented by pictures $(c)$ and $(i)$ are associated with pairs of such components ((whose pictures are obtained from each other by rotating the plane by the angle $\frac{\pi}{2})$; each of the virtual components represented by pictures $(d)$, $(e)$, $(f)$, $(g)$, and $(h)$ is associated with four real connected components of the discriminant complement, whose pictures are obtained from each other by rotating the plane by the angles $\frac{\pi}{2}$, $\pi$, and $\frac{3\pi}{2}$. The same is true for the components obtained from these nine by the operation ((18)).
Remark 19.
Picture ($c$) and its rotation by the angle $\frac{\pi}{2}$ are topologically equivalent, but the connected components of the discriminant complement realized by polynomials with these shapes are different.
Theorem 18 was conjectured in [12], and the difference of all the connected components mentioned there was proven. It remains to prove that there are no additional components, that is, the virtual components listed in Proposition 17 are associated with no more real components than indicated by this theorem.
Proposition 20.
If two generic polynomials $f$ and $\tilde{f}$ of class $\Phi_{2}$ are associated with the same virtual function, then their connected components of the space of generic polynomials of the form ((1)) are mapped to each other by an element of the group $G_{0}$.
Lemma 21.
There are exactly eight polynomials ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{2}$ that have a critical point of class $D_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{-1}{324}$. These eight polynomials form an orbit of the natural action of the group $G_{0}$.
Lemma 22.
Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by a linear transformation to the form
(19) $\pm\left[t\left(\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}\right)+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}\right]\ ,\hskip 23.49976ptt>0\ .$
In particular, the j-invariant of its principal homogeneous part is equal to 5.
Proof of Lemma 22. This lemma follows from Proposition 22 of [15] by the substitution $\tilde{x}=|a|^{\frac{-3}{4}}X$, $\tilde{y}=sign(a)|a|^{\frac{1}{4}}Y$, $t=|a|^{\frac{-5}{4}}$. $\Box$
Corollary 23.
If a polynomial of degree four in ${\mathbb{R}}^{2}$ has a critical point of class $D_{6}$ with zero critical value and a point of class $A_{3}$ with critical value $\frac{-1}{324}$, then it has the form
(20) $\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}$
in some affine coordinates.
Proof. Of all the polynomials (19), only this one has critical value $\frac{-1}{324}$. $\Box$
Corollary 24.
If a polynomial $f(x,y)$ of degree four has critical points of classes $D_{6}$ and $A_{3}$ and principal homogeneous part
(21) $x^{4}+2Ax^{2}y^{2}+y^{4}\ ,$
then $A$ is a root of the equation
(22) $\hskip 136.5733pt\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.\hskip 136.5733pt\Box$
[…]
The rest of the proof of Proposition 20 repeats the proof of Proposition 3 of [15] in § 5.4 of [15]. $\Box$
Proof of Theorem 18. Pictures ($a$) and ($b$) of Fig. 3 are realized, respectively, by the polynomials (23) and (24) with arbitrary $t>0$.
(23) $x^{4}-6x^{2}y^{2}+y^{4}+t$
(24) $x^{4}-6x^{2}y^{2}+y^{4}-2t(x^{2}+y^{2})+\frac{1}{2}t^{2}$
These polynomials are invariant under the entire group $G_{0}$. Therefore, their virtual components are associated with single real connected components of the discriminant complement.
Pictures ($c$) and ($i$) are realized respectively by polynomials (25) and (26) with arbitrary $t>0$.
(25) $(x^{2}-3t)^{2}-6(x^{2}-3t)(y^{2}-t/2)+(y^{2}-t/2)^{2}+t^{4}$
(26) $x^{4}-6x^{2}(y^{2}+t)+(y^{2}+t)^{2}-t^{2}/2.$
These polynomials are invariant under the reflections $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y),$ which generate a subgroup of index two in $G_{0}$. Thus, in each of these two cases there are no more than two connected components of the discriminant complement associated with the corresponding virtual component.
Pictures ($d$), ($e$), ($f$), ($g$), and ($h$) are realized by the equations (27), (28), (29), (30), and (31) for sufficiently small $\varepsilon>0$ and arbitrary $t>0$.
(27) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}+\varepsilon t^{3}x\right)$
(28) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}-\varepsilon t^{3}x\right)$
(29) ${\bf T}\left(x^{4}-6x^{2}y^{2}+t(x^{3}-\varepsilon^{4}t^{2}x)+(y^{2}-\varepsilon^{3}t^{2})^{2}\right)$
(30) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}-6\varepsilon txy^{2}-2\varepsilon^{4}t^{3}x+8\varepsilon^{3}t^{2}y^{2}\right)$
(31) ${\bf T}\left((x+0.7t)(x+1.3t)(x+6t)(x+7t)-6x^{2}y^{2}+y^{4}\right)$
These five polynomials are all constructed as follows. First we take the polynomial
$x^{4}-6x^{2}y^{2}+y^{4}+x^{3},$
whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at the point closest to the center of the picture. Then, we perturb this singular point by adding very small terms of lower degrees. These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [18]. All the resulting polynomials are invariant under the reflection in the axis $y=0$. Thus, in each case, the virtual component is associated with no more than four real connected components of the discriminant complement. These four components differ for the topological reasons, see [12].
For the remaining nine virtual components, the realizations of the corresponding real connected components that prove Theorem 18 in these cases can be obtained from the polynomials (23)–(31) via the operation (18). $\Box$
3. Components for $J_{10}$
3.1.
The canonical versal deformations of the singularities of the classes $J_{10}^{1}$ and $J_{10}^{3}$ consist, respectively, of polynomials
(32) $(x-Ay^{2})(x^{2}+y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}+3y^{4})y,\qquad A\in(-\infty,+\infty),$
and
(33) $(x-Ay^{2})(x^{2}-y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}-3y^{4})y,\qquad A\in(-1,1),$
with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, see [5].
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (32) and (33), respectively.
Proposition 25 (see [17]).
If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials are either the same or mapped to each other by the involution
(34) $f(x,y)\leftrightarrow f(x,-y)\ .$
In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components. $\Box$
In the following pictures, the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
3.2. Components for $J_{10}^{1}$ singularities
++-$(a)$
++-$(b)$
++-$(c)$
++-$(d)$
++-$(e)$
++-$(f)$
++-($\tilde{b}$)
++-($\tilde{c}$)
++-($\tilde{d}$)
++-($\tilde{e}$)
Figure 4. Non-discriminant perturbations of $J_{10}^{1}$ singularities
Proposition 26 (see §4.2 in [18]).
There are exactly ten topological types of zero-level sets of non-discriminant perturbations of $J_{10}^{1}$ singularities, see Fig. 4. $\Box$
Proposition 27 (see [12]).
There are exactly ten virtual components of the formal graph of $J_{10}^{1}$ singularities. These virtual components are represented by polynomials, the topological shapes of whose zero-level sets are shown on Fig. 4. $\Box$
Theorem 28.
There are exactly thirteen connected components of the discriminant complement in the space of all polynomials of class $\Omega_{1}$. Each of the virtual components shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig. 4 is associated with a single real connected component. Each of the virtual components/topological types shown in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((34)).
Proof of Theorem 28. The topological types shown in pictures ($a$), ($b$), ($c$), and ($d$) are represented respectively by the following polynomials:
$x(x^{2}+y^{4}+t),$
$(x-2t)((x+t)^{2}+y^{4}-t^{2}),$
$(x-2t)((x+t)^{2}+y^{4}-3ty^{2}+2t^{2}),$
$(x-y^{2}+2t)(x^{2}-2tx+y^{4}-5t^{2})-t^{3}/10$
with arbitrary $t>0$. These polynomials are all invariant under the involution (34). Therefore, the corresponding virtual components are associated with single connected components of the discriminant complement.
The realizations of pictures (b~)($\tilde${b}), (c~)($\tilde${c}), and (d~)($\tilde${d}) can be obtained from the realizations of pictures $(b),$ $(c)$, and $(d)$ by the involution
(35) $f(x,y)\mapsto-f(-x,y).$
All of them are also invariant under the involution (34).
No polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(f)$ can belong to the same connected component as its image f(x,y)f(x,-y) under the involution (34). Indeed, the polynomials (32) are of degree three in the variable $x$. Therefore, any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times.
An invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ realizing the picture $(e)$ from its mirror image f(x,y)f(x,-y) is as follows. Four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take three arbitrary points inside the rightmost three ovals. There is a single parabola of the form $x=ay^{2}+by+c$ that passes through these three points. The restriction of the polynomial $f$ to this parabola is a polynomial of degree at most six in the coordinate $y$. This restriction has six roots, so it is of degree exactly six. This parabola cannot intersect the fourth oval, hence it passes either above or below it with respect to the $x$ coordinate. The position of the parabola relative to the fourth oval is independent of the choice of three points inside the ovals, and remains unchanged along the connected components of the discriminant complement. However, these relative positions are different for any polynomial with the topological shape of the zero-level set shown in picture $(e)$ and for its mirror image f(x,y)f(x,-y). The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
3.3. Components for $J_{10}^{3}$ singularities
---++++++$(a)$
---++++++$(b)$
---++++++$(c)$
---++++++$(d)$
---++++++$(e)$
---++++++$(f)$
---++++++$(g)$
---++++++$(h)$
---++++++$(i)$
---++++++$(j)$
---++++++$(k)$
---++++++$(l)$
---++++++$(m)$
Figure 5. Perturbations of $J_{10}^{3}$
Proposition 29 (see [12]).
There are exactly 23 virtual components of class $\Omega_{3}$. The shapes of the zero-level sets of polynomials $f(x,y)$ representing some thirteen of them are shown in Fig. 5 (see also Fig. 29 in [18]). In all these cases except those shown in pictures $(h)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((35)), represent different virtual components.
We will denote by (a~)($\tilde${a}), (b~)($\tilde${b}),…the pictures obtained from the corresponding pictures $(a)$, (b)(b),…by the involution induced by the transformation (35), that is, by the reflection in a horizontal line and simultaneous change of all signs.
Remark 30.
Pictures $(d)$ and $(e)$ of Fig. 5 are topologically equivalent, however it follows from the Bezout’s theorem that they represent different connected components.
Theorem 31.
There are exactly 33 connected components of the complement of the discriminant variety of the versal deformation ((33)). Each virtual component represented by a picture of Fig. 5 in the list
(36) (h),(h~),(i),(i~),(j),(j~),(k),(k~),(g),(l)(h),\ ($\tilde${h}),\ (i),\ ($\tilde${i}),\ (j),\ ($\tilde${j}),\ (k),\ ($\tilde${k}),\ (g),\ (l)
[…]
Proof. The difference of all these 33 connected components easily follows from the topological considerations and Bezout’s theorem, see [12]. By Proposition 25, it remains to be shown that, in all cases not covered by the list (36), the involution (34) does not change the real connected component.
The polynomials
$(x-t)^{3}-(x-t)y^{4}+3(x-t)^{2}+5(x-t)y^{2}-4(x-t)-Ct^{3}$
with arbitrary $t>0$ realizes the picture $(a)$ for $C<6-\frac{14}{3}\sqrt{\frac{7}{3}}$, the picture $(b)$ for $C\in\left(6-\frac{14}{3}\sqrt{\frac{7}{3}},-\frac{1}{2}\right)$, and the picture $(d)$ for $C\in\left(-\frac{1}{2},0\right)$. All of these polynomials are even with respect to the variable $y$, so their virtual components are represented by single connected components of the complements of the discriminants.
Picture $(m)$ is represented by the polynomials
$x(x-y^{2}-t)(x+y^{2}+t),\qquad t>0,$
which are also self-symmetric.
The polynomials
$(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+3t(x-t)y^{2}+Ct^{3}$
realize picture $(c)$ for any $C\in(0,1/2))$. They are also invariant under the involution (34).
The polynomials
$(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+5t(x-t)y^{2}-4t^{2}(x-t)+\frac{t^{2}}{10}y^{2}-0.35t^{3}$
realize picture $(f)$.
Figure 6.
To realize the picture $(e)$, we start with the polynomial
(37) $x(x+(y-2)^{2}-2)(x+(y+2)^{2}-2)$
whose zero-level set is shown in Fig. 6. Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal quasihomogeneous part will be of class $J_{10}^{3}$. The polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values. It will also have two critical points in the complex domain. Then we add a very small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain. The topological picture of the zero-level set of the resulting polynomial will be the desired one. Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal quasihomogeneous part of the resulting polynomial will have the canonical form, see Table 2. Finally, we apply the operator $\bf T$} to reduce the polynomial to the form (1). The resulting polynomial realizes the picture $(e)$, is of class $\Omega_{3}$, and is invariant under the involution (34). $\Box$
4. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at the $X_{9}$ and $J_{10}$ singularities proves all of the conjectures in [11] concerning the number of local Petrovskii lacunas at the corresponding singular points of wavefronts.
In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity when the dimension $N$ and the index $i_{+}$ are both even is component $(a)$ of Fig. 1. The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure. The four lacunas at the $X_{9}^{2}$ singularity, with even $N$ and odd $i_{+}$, are shown in picture $(i)$ of Fig. 3, or in its rotation by the angle $\frac{\pi}{2}$, or in the pictures obtained by applying the operation (18) to these two. The unique lacuna at the $J_{10}^{3}$ singularity, with even $N$ and odd $i_{+}$, is shown in picture $(m)$ of Fig. 5.