### v2

# Complements of discriminants of real parabolic function singularities. II

Abstract.
We provide a complete list of the connected components of the spaces of non-discriminant functions within standard versal deformations of function singularities of classes $X_{9}$, $J_{10}$ and $P_{8}^{1}$ (as well as a partial list for the remaining class, $P_{8}^{2}$). Thus, we prove (and improve in one particular case) the corresponding conjectures from the previous work [18] with the same title. As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs. In particular, we discover a new local lacuna at the $P_{8}^{2}$ singularities. We also show that the complements of the discriminant varieties of $X_{9}^{+}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, unlike all simple singularities.
Key words and phrases:
discriminant, singularity, surgery, versal deformation, isotopy classification
1991 Mathematics Subject Classification:
Primary: 14Q30. Secondary: 14B07, 14P25
1. Introduction
$1.$
Discriminant varieties of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of these functions are singular. These varieties appear in many problems of PDE theory and integral geometry, often under the name “wavefronts”, see e.g. [1], [2], [3], [8], [15][16], [17].
Many important functions in these theories arise as integral transforms. These functions are defined on the parameter spaces of certain function deformations (called generating families) and are equal to the integrals along the cycles in the zero-level sets of the corresponding functions. These integrals are regular outside the discriminant varieties of the deformations. Their analytic behavior near the discriminants depends on the singularity type of the deformed discriminant function and on the connected component of the complement of the discriminant. Thus, enumerating these components is one of the first steps in studying the analytical properties of these integral transforms.
Definition 1 (see [2]).
A function singularity is a germ of a regular analytic function $({\mathbb{R}}^{n},0)\to({\mathbb{R}},0)$ at its critical point. Two function singularities at points $a$ and $b$ are equivalent if they can be transformed to each other via a local diffeomorphism $({\mathbb{R}}^{n},a)\to({\mathbb{R}}^{n},b)$. Two function singularities, $f(x_{1},\dots,x_{n})$ and $\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})$, are stably equivalent if they become equivalent after the summation with non-degenerate quadratic forms in additional variables, that is,
$f(x_{1},\dots,x_{n})+Q_{2}(y_{1},\dots,y_{N-k})\sim\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})+\tilde{Q}_{2}(\tilde{y}_{1},\dots,\tilde{y}_{N-m}).$
An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin in ${\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic. The discriminant of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a real critical point with zero critical value (or, equivalently, the zero-level set $f_{\lambda}^{-1}(0)$ is singular near real space).
For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [2]) deformation. The discriminant sets of all versal deformations of stably equivalent singularities are topologically equivalent, and the components of their complements are in a one-to-one correspondence. Therefore, for any stable equivalence class of singularities it is sufficient to enumerate the components for only one versal deformation of one function of this class.
The list of equivalence classes of function singularities begins with simple singularities, which were distinguished by V. Arnold, see [2]. All of these singularities are stably equivalent to functions in two variables, and the entire geometric study of their deformations and discriminants reduces to the study of functions in two variables. The components of discriminant complements for these singularities were listed earlier, see § 1.3.
Table 1. Normal forms of simple, $X_{9}$ and $J_{10}$ singularities in two variables and $P_{8}$ singularities in three variables
Notation Normal form Restriction
$A_{2k-1}$ $\pm x^{2k}\pm y^{2}$ $k\geq 1$
$A_{2k}$ $x^{2k+1}\pm y^{2}$ $k\geq 1$
$D_{2k}^{\pm}$ $x^{2}y\/\pm y^{2k-1}$ $k\geq 2$
$D_{2k-1}$ $\pm(x^{2}y\/+y^{2k-2})$ $k\geq 3$
$E_{6}^{\pm}$ $x^{3}\pm y^{4}$
$E_{7}$ $x^{3}+xy^{3}$
$E_{8}$ $x^{3}+y^{5}$
$X_{9}^{\pm+}$ $\pm(x^{4}+2Ax^{2}y^{2}+y^{4})$ $|A|<1$
$X_{9}^{1}$ $xy(x^{2}+2Axy+y^{2})$ $|A|<1$
$X_{9}^{2}$ $x^{4}+2Ax^{2}y^{2}+y^{4}$ A<1A<-1
$J_{10}^{1}$ $(x^{2}+y^{2})(x-\gamma y^{2})$ $\gamma\in(-\infty,\infty)$
$J_{10}^{3}$ $(x^{2}-y^{2})(x-\gamma y^{2})$ $\gamma\in(-1,1)$
$P_{8}^{1}$ $x^{3}+y^{3}+z^{3}-3Axyz$ $A<1$
$P_{8}^{2}$ $x^{3}+y^{3}+z^{3}-3Axyz$ $A>1$
In this work we list analogous components for the next natural class in the hierarchy of function singularities, the parabolic singularities, which were also distinguished by V. Arnold. The classification of complex parabolic singularities up to the stable equivalence consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$. The classification of their real forms includes more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$. Functions of classes $X_{9}$ and $J_{10}$ are stably equivalent to the functions in two variables, while the $P_{8}$ singularities are stably equivalent to the non-degenerate cubic forms in three variables. Any function singularity in two variables of class $X_{9}^{*}$ or $J_{10}^{*}$, as well as any function in three variables of one of classes $P_{8}^{*}$ can be reduced to a normal form indicated in Table 1 by an appropriate choice of local coordinates centered at the critical point. Additionally, this table provides the normal forms of simple singularities $A_{k},D_{k},$ and $E_{k}$, into which the parabolic singularities can split after small perturbations.
We give complete lists of components for all parabolic singularity classes except $P_{8}^{2}$, for which we have only partial results.
Table 2. Numbers of rigid isotopy and topological classes of non-discriminant perturbations of parabolic singularities
Singularity class $X_{9}^{\pm}$ $X_{9}^{1}$ $X_{9}^{2}$ $J_{10}^{1}$ $J_{10}^{3}$ $P_{8}^{1}$ $P_{8}^{2}$
Isotopy 7 14 52 13 33 7 $\geq 15$
Topology 6 12 50 10 31
O. Viro [24] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities. By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood. The topological type is an invariant of the components of the discriminant complements. The “rigid isotopy” classification of parabolic singularity perturbations given in the present work is only slightly stricter than the topological classification, see Table 2. Counting the components of the discriminant complements of $P_{8}$ singularities is related to (though is different from) the isotopy classification of nonsingular cubic surfaces in ${\mathbb{R}}^{3}$, which is described in detail in [7]. This classification consists of five and ten classes for $P_{8}^{1}$ and $P_{8}^{2}$ singularities, respectively.
1.2. On the method
In [18], a system of invariants of connected components of the complements of the discriminant varieties of function singularities is described. Complete lists of possible values of these invariants for parabolic singularities are given in [18], and all of these values are realized by concrete polynomials. In the same paper, we also conjectured that the resulting lists of connected components are complete. These lists consist of 7, 15, 7, 14, 52, 11, and 33 components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [18] consist of pairs of different connected components. Also, we present below partial results for the case $P_{8}^{2}$, which prove that at least 10 out of 15 classes listed in [18] are indeed represented by only one connected component each. Thus, the number 11 above is replaced in Table 2 by 13, and the number 15 is still under suspicion. These results are summarized in Theorems 5, 13, 18, 28, 31, 37, and 41 below.
1.3. On the discriminant complements for simple singularities
E. Looijenga [13] implicitly listed the local connected components of the complements of the discriminants of simple singularities. O. Viro [24] listed all possible topological types of non-discriminant functions representing these components. In [19], it was shown that each topological type on this list corresponds to exactly one connected component of the discriminant complement. For simple singularities with Milnor numbers at most six this fact also follows from the results of [14]. Unfortunately, I only read the work [24] containing the enumeration of topological types of simple and parabolic function perturbations in two variables after the publication of [18] and [19].
1.4. 1-cohomology of discriminant complements
E. Looijenga [13] proved that all local components of discriminant complements of real simple singularities are topologically trivial. Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 12 and 39.
1.5. An application to hyperbolic PDEs.
In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [1], [8], [16]). Near the non-singular points of wavefronts (corresponding to the Morse functions) the lacunas were enumerated in [6] and [5]. Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [8], and near all simple singularities in [15]. In [17], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that the list is complete. We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5. For the singularities of classes $P_{8}^{1}$, $X_{9}^{*}$ and $J_{10}^{*}$, these results follow directly from our complete lists of components of discriminant complements. Our partial results on enumerating components for the $P_{8}^{2}$ singularities are sufficient to prove that our new list of local lacunas is also complete in that case.
2. Results for the $X_{9}^{*}$ classes
2.1.
Real parabolic function singularities of type $X_{9}$ form four stable equivalence classes $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$.
The standard versal deformations of functions in all these classes have the same form: they are families of polynomials
(1) $f_{A}(x)+\lambda_{1}+\lambda_{2}x+\lambda_{3}y+\lambda_{4}x^{2}+\lambda_{5}xy+\lambda_{6}y^{2}+\lambda_{7}x^{2}y+\lambda_{8}xy^{2},$
depending on nine parameters $\lambda_{1},\dots,\lambda_{8},$ and $A$, where $f_{A}$ is the corresponding normal form from Table 1. According to versality theory (see, for example, [2]), studying all other deformations of singularities of stable equivalence classes $X_{9}^{*}$ can be reduced to studying these families.
Definition 2.
A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (1) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ consists of four different lines in ${\mathbb{C}}^{2}$. Denote by $\Phi_{+}$, $\Phi_{1}$, and $\Phi_{2}$ the spaces of all polynomials of the form (1) with principal parts of classes $X_{9}^{+}$, $X_{9}^{1}$, and $X_{9}^{2}$, respectively.
The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$. The discriminant subvarieties of these spaces consist of functions that have at least one real critical point with zero critical value. These varieties divide the spaces $\Phi_{*}$ into connected components. Enumerating of these components is our goal in this section.
2.2. Virtual functions and virtual components
In §2 of [18] we introduced the set-valued invariant of connected components of the discriminant complements. Let us briefly review this construction in the case of $X_{9}$ singularities.
The virtual function associated with a generic real polynomial is a collection of topological characteristics of the complex zero-level set of this polynomial, depending also on the collections of paths in ${\mathbb{C}}^{1}$ defining the bases of vanishing cycles in the homology of this level set. Specifically, the data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
  • the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented vanishing cycles forming the basis of the two-dimensional homology group of the submanifold of ${\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
  • the string of intersection indices of these nine cycles with the set of real points,
  • the Morse indices of real critical points of $f$, and
  • the numbers of positive and negative critical values of $f$.
The (abstract) virtual functions of class $\Phi_{*}$ include all virtual functions associated with generic real polynomials of this class, as well as all analogous collections of data that can be obtained from them via standard flips that model surgeries of real functions. The formal graph of any type $X_{9}^{*}$ is the graph, whose vertices correspond to all the virtual functions of the corresponding class $\Phi_{*}$, and whose edges correspond to the standard flips. The virtual components of the formal graph are the connected subgraphs, into which the formal graph splits when all edges modeling the surgeries that cross the discriminant are removed. The set-valued invariant of a generic polynomial $f$ of class $\Phi_{*}$ is the virtual component of the formal graph containing the virtual functions associated with this polynomial. This virtual component is denoted by $S(f)$.
These objects have the following obvious property.
Proposition 1.
If two generic polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$, S(f~)S($\tilde${f}) are the same. $\Box$
The formal graphs of all simple and parabolic singularities are finite.
Enumerating the virtual components of such a graph is a combinatorial problem that can be solved using a computer program. The work [18] provides the results of these computations for all parabolic singularities, as well as the immediate consequences of these computations in the terms of real functions. The remaining problem is to find all real connected components with this value of the set-valued invariant for each virtual component.
The results of [10] imply that in the case of parabolic singularities all standard flips of virtual functions can be realized by real surgeries of generic real functions. In particular, the following statement holds.
Proposition 2 (see [21], Proposition 3).
For any class of polynomials $\Phi_{+}$, $\Phi_{1}$, or $\Phi_{2}$,
a)) every virtual function of this class is associated with a generic real polynomial of the same class;
b)) every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial $\tilde{f}$ from the same connected component of the complement of the discriminant as $f$. $\Box$
Remark 3.
In the statement of Proposition 3 of [21], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with nondegenerate principal homogeneous parts. However, the entire construction proving this proposition in [21] can be performed within the spaces of the versal deformations (1). Thus, it also proves the above Proposition 2.
These constructions enable us to estimate from above the number of isotopy classes of generic Morse perturbations associated with any given virtual function. These estimates are given in Propositions 8, 15, and 20 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately. Consequently, we obtain estimates of the number of connected components of the complement of a real discriminant which are associated with each virtual component. Specifically, in the cases $X_{9}^{+}$ and $X_{9}^{2}$, all real connected components of discriminant complements associated with the same virtual component can be obtained from each other by the action of a transformation group of order eight. In the case $X_{9}^{1}$ the analogous group is of order four. Then, for each virtual component in the list from [18], we compare all real components associated with this virtual component and determine which of these components are identical. This comparison provides the results of the present work concerning the $X_{9}$ singularities. The strategy for the $J_{10}$ and $P_{8}$ singularities is exactly the same.
Remark 4.
In [21], [22], and [23], we use the same strategy with a different definition of virtual components. In these works, we classify the Morse perturbations of function singularities up to isotopy within the space of Morse functions instead of the space of non-discriminant functions. Thus, the virtual components of the formal graph are determined there by removing the edges that change the Morse type.
Definition 3.
For any Morse polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$, the number $\mbox{Ind}(f)$ equals the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
It is easy to see that this number is an invariant of the connected components of the discriminant complements (and of associated virtual components).
2.3. Enumeration of components for $X_{9}^{+}$ singularities
In this subsection, we prove the following statement, conjectured in [18].
++$(a)$
(b)(b)++
(c)(c)++
++$(d)$
(e)(e)++
++$(f)$
++$(g)$
Figure 1. Perturbations of $X_{9}$ singularity
Theorem 5.
There are exactly seven connected components in the space of non-discriminant functions of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials representing these components are shown in Fig. 1.
Remark 6.
Pictures $(f)$ and $(g)$ of Fig. 1 are topologically equivalent. However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$. Figure 1 coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves. In our case, however, the isotopies are allowed only within a much smaller space $\Phi_{+}$. In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
The starting point of the proof of Theorem 5 is the following proposition.
Proposition 7 (see [18]).
There are exactly seven virtual components of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials associated with virtual functions representing these virtual components are shown in Fig. 1. $\Box$
Definition 4.
Denote by $G_{0}$ the group of linear transformations of ${\mathbb{R}}^{2}$ generated by reflections $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(y,x)$
Clearly, this group is isomorphic to the group of symmetries of the square, in particular is of order eight. This group acts on the space $\Phi_{+}$ (each transformation $\Gamma:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ moves the polynomial $f$ to $f\circ\Gamma$). This action preserves the set of generic polynomials.
Proposition 8.
If two generic polynomials, $f$ and $\tilde{f}$, of class $\Phi_{+}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of class $\Phi_{+}$ containing $f$ and $\tilde{f}$ are mapped to each other by an element of the group $G_{0}$.
Definition 5.
For any polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree four with principal homogeneous part $x^{4}+2Ax^{2}y^{2}+y^{4}$, denote by ${\bf T}(f)$ the unique polynomial of the form (1) with the same principal part, which is obtained from $f$ by a parallel translation of ${\mathbb{R}}^{2}$. (This is the translation by the vector $\frac{1}{4}(a,b),$ where $a$ and $b$ are the coefficients of $f$ at the monomials $x^{3}$ and $y^{3}$).
Lemma 9 (see [21], Lemma 4).
A polynomial of the form ((1)) with a positive principal homogeneous part has a critical point of class $E_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{27}{256}$ if and only if it is one of the four polynomials
(2) $\hskip 76.82234pt{\bf T}(x^{4}\pm x^{3}+y^{4})\quad\mbox{ or }\quad{\bf T}(x^{4}+y^{4}\pm y^{3})\ .\hskip 76.82234pt\Box$
Proof of Proposition 8. This proof almost coincides with the proof of Proposition 3 in §3.5 of [21]. In the next paragraph, we present the outline of this proof; we refer to [21] for all common details.
Using the Lyashko–Looijenga covering (in the form described in [10]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of not strictly Morse functions). Since the virtual function of $\tilde{f}$ is the same as of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. (For this existence statement, the parabolicity of $X+9$ singularities is essential.) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map: this is the map induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$. Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are the same, or they are mapped into each other by this automorphism. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads us to the situation considered previously. $\Box$
Proof of Theorem 5. By Proposition 8, any virtual component mentioned in Proposition 7 is associated with at most eight real connected components. Let us prove that in fact each virtual component represented by a polynomial shown in Fig. 1 is associated with a single real connected component.
Pictures $(a)$, $(c)$, $(b)$, $(f)$ of Fig. 1 are realized, respectively, by polynomials
$x^{4}+y^{4}+t,$
$x^{4}+y^{4}-t,$
$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(0,1/4),$
$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\quad\alpha\in(1/4,1/2),$
with arbitrary $t>0$. All of these polynomials are invariant under the entire group $G_{0}$.
Picture (d) can be represented by the polynomial
(3) $x^{4}+2Ax^{2}y^{2}+y^{4}+2x^{2}-2y^{2}+1/2$
with arbitrary $A\in(-1,1)$. This polynomial is invariant under the transformations $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y)$. Thus, its $G_{0}$-orbit consists only of itself and the polynomial
(4) $x^{4}+2Ax^{2}y^{2}+y^{4}-2x^{2}+2y^{2}+1/2\ .$
The polynomials (3) and (4) are connected by the family of polynomials
(5) $x^{4}+2Ax^{2}y^{2}+y^{4}+2(x\cos\varphi+y\sin\varphi)^{2}-2(x\sin\varphi-y\cos\varphi)^{2}+1/2,$
that depend on the parameter $\varphi\in[0,\pi/2]$.
Lemma 10.
[…]
Proof. If we replace $A$ by 1 in (5), then all these polynomials will be obtained from each other by the rotations of ${\mathbb{R}}^{2}$ by angles $\varphi\in[0,\pi/2]$. Each of these polynomials has only three real Morse critical points: one at the origin and two on the unit circle. If $A=1-\varepsilon$ with sufficiently small $\varepsilon$, then the real zero-level set deforms only slightly in the domain $\{x^{2}+y^{2}\leq 2\}.$ In particular, it maintains its topological structure in this domain for any $\varphi$. It also remains empty outside this domain. $\Box$
Thus, the connected component of the space of non-discriminant polynomials containing (3) is also invariant under the transformation $(x,y)\mapsto(y,x)$.
The topological type of the curve shown in picture $(e)$ can be realized by any of the four polynomials
(6) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(x^{3}-3xy^{2})+\varkappa),$
(7) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(y^{3}-3x^{2}y)+\varkappa)$
with arbitrary $A\in(-1,1)$ and sufficiently small $\varkappa>0$ (where the restriction on $\varkappa$ depends on $A$). These four polynomials form an orbit of the group $G_{0}$. All of them belong to the one-parametric family of polynomials
(8) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}+\mbox{Re}(e^{i\tau}(x+iy)^{3}))+\varkappa),$
$\tau\in[0,2\pi]$. For $A=1$, all these polynomials are obtained from each other by the rotations of ${\mathbb{R}}^{2}$. They are bounded from below, have a critical point of class $D_{4}^{-}$ with critical value $\varkappa$ at the origin, three real minimum points with critical value $\frac{27}{256}+\varkappa$, and no other critical points. Therefore, for $\varkappa\in\left(0,\frac{27}{256}\right)$ all these polynomials realize the topological picture $(e)$. Analogously to the proof of Lemma 10, taking $A<1$ very close to 1 will not change these topological pictures for any $\tau$. So, the connected component of the discriminant complement containing these polynomials is invariant under the action of the group $G_{0}$.
The picture $(g)$ can be realized by either of the four polynomials
(9) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10x(x^{2}-3y^{2})+15(x^{2}+y^{2})-5)\ ,$
(10) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10y(y^{2}-3x^{2})+15(x^{2}+y^{2})-5)$
with sufficiently small $\varepsilon>0$. These four polynomials form an orbit of the group $G_{0}$.
The family of polynomials
(11) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}+10\mbox{Re}\left(e^{i\tau}(x+iy)^{3}\right)+15(x^{2}+y^{2})-5),$
$\tau\in[0,2\pi],$ contains all four polynomials (9), (10) for $\tau=0,$ $\frac{\pi}{2}$, $\pi$, and $\frac{3\pi}{2}$. All elements of this family lie in the same connected component of the complement of the discriminant. This completes the proof of Theorem 5. $\Box$
Remark 11.
To realize the pictures ($d$), ($e$), and ($g$) by arbitrarily small perturbations of the $X_{9}^{+}$ singularities, we can multiply by $t^{4-k}$ all monomials of degree $k$ in the formulas (3), (6), and (9).
2.4. The 1-cohomology of discriminant complements of $X_{9}^{+}$ singularities
Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, that is, the space of $k$-point subsets of the plane. For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$. The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$. Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are $k\geq 2$ non-nested ovals. Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(U)$ is defined as follows. For any polynomial $f\in U,$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals. These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $U$. The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $U\to B({\mathbb{R}}^{2},k)$. Clearly, these maps defined by different cross-sections are homotopic and define the same homomorphism of cohomology rings.
Define the class $W\in H^{1}(U)$ as the class obtained by this map from the generator of the group $H^{1}(B({\mathbb{R}}^{2},k))$.
Theorem 12 (cf. [20], § 1.8.3).
The class $W$ takes unit values on some 1-cycles in each of the connected components shown in pictures ((d)), ((e)), and ((g)) of Fig. 1.
Proof. These cycles are the families (5), (8), and (11). $\Box$
2.5. The components for the $X_{9}^{1}$ singularities
Define the group $G_{1}$ as the group of linear transformations of ${\mathbb{R}}^{2}$ generated by the symmetries $(x,y)\mapsto(y,x)$ and $(x,y)\mapsto(-x,-y)$. Clearly, this group is isomorphic to ${\mathbb{Z}}_{2}^{2}$. It acts on the space $\Phi_{1}$ of polynomials (1) with principal part
(12) $xy(x^{2}+2Axy+y^{2}),\quad|A|<1,$
and preserves the set of generic polynomials. Polynomials obtained from each other by this action have equal associated virtual functions and so are not separated by the set-valued invariant.
Additionally, the involution
(13) $f(x,y)\mapsto-f(x,-y)$
acts on the space $\Phi_{1}$. It preserves the set of discriminant polynomials and hence permutes the connected components of its complement.
This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y). These involutions have no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components (real or virtual) related by the involution is always equal to 1-1.
For any polynomial $f$ with a principal homogeneous part of the form (12), its zero-level set goes to the infinity along four branches: two vertical and two horizontal. If the polynomial is non-discriminant, then each vertical branch is matched with the horizontal branch belonging to the same component of the zero-level set. The involution (13) acts on the zero-level sets as the reflection in the line $\{y=0\}$. In particular, it always alters the matching of the vertical and horizontal branches.
++--++$(a)$
++--++$(b)$
++--++$(c)$
++--++$(d)$
++--++$(e)$
++--++$(f)$
++--++$(g)$
Figure 2. Perturbations of $X_{9}^{1}$ singularity
Theorem 13.
There are exactly 14 connected components of the complement of the discriminant in the space of polynomials ((1)) of class $\Phi_{1}$. The shapes of the zero-level sets of polynomials representing some seven of them are shown in Fig. 2, and the remaining seven can be obtained from these by the involution ((13)).
The proof of this theorem is based on the following two propositions.
Proposition 14 (see [18]).
There are exactly ten virtual components of type $X_{9}^{1}$. They split into five pairs of virtual components related to each other by the involution induced by ((13)). The topological types of zero-level sets of polynomials associated with elements of some five of the virtual components ((one from each pair)) are shown in pictures ((a)), ((b)), ((c)), ((d), and ((f)) of Fig. 2. $\Box$
Proposition 15.
If two generic polynomials $f$ and $\tilde{f}$ of class $\Phi_{1}$ are associated with the same virtual function, then their connected components in the space of generic polynomials are moved to each other by an element of the group $G_{1}$.
Definition 6.
For any polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree four with the principal homogeneous part (12), denote by ${\bf\tilde{T}}(f)$ the unique polynomial of the form (1) with the same principal part, which is obtained from $f$ by a parallel translation of the plane. (This is the translation by the vector $(\alpha,\beta),$ where $\alpha$ and $\beta$ are the coefficients of $f$ at the monomials $y^{3}$ and $x^{3}$).
Lemma 16.
A polynomial of the form ((1)) with the principal part ((12)) has a critical point of class $E_{7}$ with zero critical value and a critical point of class $A_{2}$ with critical value $3^{\frac{15}{2}}2^{-8}$ if and only if it is one of four polynomials
(14) ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm x^{3})\ \mbox{ \rm or }\ {\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm y^{3})\ .$
Proof. Let $f$ be a polynomial that satisfies the conditions of this lemma. By Lemma 5 of [21], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (12). In the affine coordinates $(x,y)$ centered at the $E_{7}$ point, the monomials of degree $\leq 2$ of this polynomial vanish, and the homogeneous part of degree three has the form $\alpha x^{3}$ or $\alpha y^{3}$, $\alpha\neq 0$. The critical value of this function at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$. This value is equal to $3^{\frac{15}{2}}2^{-8}$ if and only if this sign is ++ and $\alpha$ equals $1$ or 1-1. $\Box$
Proof of Proposition 15. The deduction of this proof from Lemma 16 essentially repeats the proof of Proposition 8 with the following simplification. In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial involutions that commute with the Lyashko–Looijenga covering (see [21], § 4.5). Therefore, the statement of Proposition 15 follows immediately from the fact that all four polynomials (14) belong to the same orbit of the group $G_{1}$. $\Box$
Proof of Theorem 13. Proposition 15 implies that any virtual component of the formal graph of type $X_{9}^{1}$ is associated with at most four real connected components mapped to each other by the group $G_{1}$. Below we prove that six of the virtual components described in Proposition 14 are associated with single real connected component each, while each of the four remaining virtual components is associated with a pair of real connected components that are mapped into each other by the central symmetry in ${\mathbb{R}}^{2}$.
Pictures $(a)$, $(b)$, and $(c)$ of Fig. 2 are realized by the polynomials
$xy(x^{2}+y^{2})-t,$
$xy(x^{2}+y^{2})-t(x^{2}+y^{2})+\alpha t^{2},\quad\alpha\in(0,1/2),$
and
$xy(x^{2}+y^{2}-t)-\alpha t^{2},\quad\alpha\in(0,1/8)$
for any $t>0$. All of these polynomials are invariant under the action of the group $G_{1}$. The polynomials obtained from them by the involution (13) and representing other three connected components of the discriminant complement are also invariant under this action.
Picture $(d)$ (respectively, $(e)$) is realized by the polynomial
(15) ${\bf\tilde{T}}\left((x^{2}+y^{2}-t^{2})(x+0.9t)(y+0.9t)+\alpha t^{4}\right),$
$\alpha\in(0,0.003)$, with positive (respectively, negative) $t$. These two pictures are topologically equivalent. However, it easily follows from Bezout’s theorem that these polynomials represent different connected components of the discriminant complement in the space $\Phi_{1},$ see [21].
Each polynomial (15) is invariant under the transformation $(x,y)\mapsto(y,x)$. The transformation $(x,y)\mapsto(-x,-y)$ maps each such polynomial to the analogous polynomial with the opposite value of parameter $t$. In particular, it permutes the connected components described by pictures $(d)$ and $(e)$.
Pictures $(f)$ and $(g)$ of Fig. 2 are realized by polynomials
(16) $xy\left((x-2t)^{2}+(y-2t)^{2}-t^{2}\right)-t^{4}=0$
with positive and negative values of $t$, respectively. Each of these polynomials is invariant under the transformation $(x,y)\mapsto(y,x)$. The transformation $(x,y)\mapsto(-x,-y)$ maps each polynomial (16) to the polynomial with the opposite value of the parameter $t$, thus it permutes two connected components represented by these polynomials.
The same considerations hold for components obtained from these by the involution (13). This concludes the proof of Theorem 13. $\Box$
2.6. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way. All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
In addition, the operation
(17) $f(x,y)\mapsto-f\left(\frac{x+y}{\sqrt{2}},\frac{x-y}{\sqrt{2}}\right)$
acts on the space $\Phi_{2}$. This operation always alters the real connected components of the discriminant complement, as well as the associated virtual components of type $\Phi_{2}$. Indeed, it maps any generic polynomial with value $i$ of the invariant Ind to a polynomial with value 3i-3-i. This operation is of order eight. All even iterations of it act as rotations by angles of $\frac{\pi k}{2}$ and thus belong to the group $G_{0}$. So, the entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation.
----++$(a)$
----++$(b)$
----++$(c)$
----++++$(d)$
----++++$(e)$
----++++$(f)$
----++++$(g)$
----++++$(h)$
++++++--$(i)$
Figure 3. Non-discriminant perturbations of $X_{9}^{2}$
Proposition 17 (see [18]).
There are exactly 18 virtual components of class $\Phi_{2}$. The shapes of the zero-level sets of the polynomials representing nine of these virtual components are shown in Fig. 3 ((see also Fig. 33 in [24])), and the polynomials representing the remaining nine virtual components are obtained from these by the operation ((17)).
Theorem 18.
There are exactly 52 connected components of the discriminant complement in the space $\Phi_{2}$.
Specifically, the virtual components represented by pictures $(a)$ and $(b)$ of Fig. 3 are associated with single real connected components of the discriminant complement; the virtual components represented by pictures $(c)$ and $(i)$ are associated with pairs of such components ((whose pictures are obtained from each other by rotating the plane by the angle $\frac{\pi}{2})$; each of the virtual components represented by pictures $(d)$, $(e)$, $(f)$, $(g)$, and $(h)$ is associated with four real connected components of the discriminant complement, whose pictures are obtained from each other by rotating the plane by the angles $\frac{\pi}{2}$, $\pi$, and $\frac{3\pi}{2}$. The same is true for the components obtained from these nine by the operation ((17)).
Remark 19.
Picture ($c$) and its rotation by the angle $\frac{\pi}{2}$ are topologically equivalent, but the connected components of the discriminant complement realized by polynomials with these shapes are different.
Theorem 18 was conjectured in [18], and the difference of all the connected components mentioned there was proven. It remains to prove that there are no additional components, that is, the virtual components listed in Proposition 17 are associated with no more real components than indicated in this theorem.
Proposition 20.
If two generic polynomials $f$ and $\tilde{f}$ of class $\Phi_{2}$ are associated with the same virtual function, then their connected components of the space of generic polynomials of the form ((1)) are mapped to each other by an element of the group $G_{0}$.
Lemma 21.
There are exactly eight polynomials ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{2}$ that have a critical point of class $D_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{-1}{324}$. These eight polynomials form an orbit of the natural action of the group $G_{0}$.
Lemma 22.
Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by an affine transformation of ${\mathbb{R}}^{2}$ to the form
(18) $\pm\left(t\left(\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}\right)+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}\right)\ ,\quad t>0\ .$
In particular, the j-invariant of its principal homogeneous part is equal to 5.
Proof of Lemma 22. This lemma follows from Proposition 22 of [21] by the substitution $\tilde{x}=|a|^{\frac{-3}{4}}X$, $\tilde{y}=sign(a)|a|^{\frac{1}{4}}Y$, $t=|a|^{\frac{-5}{4}}$. $\Box$
Corollary 23.
If a polynomial of degree four in ${\mathbb{R}}^{2}$ has a critical point of class $D_{6}$ with zero critical value and a point of class $A_{3}$ with critical value $\frac{-1}{324}$, then it has the form
(19) $\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}$
in some affine coordinates.
Proof. Of all the polynomials (18), only the polynomial (19) has critical value $\frac{-1}{324}$. $\Box$
Corollary 24.
If a polynomial $f(x,y)$ of degree four with principal homogeneous part
(20) $x^{4}+2Ax^{2}y^{2}+y^{4}$
has critical points of classes $D_{6}$ and $A_{3}$, then $A$ is a root of the equation
(21) $\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.$
Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [10].
[…]
The rest of the proof of Proposition 20 repeats the proof of Proposition 3 of [21] in § 5.4 of [21]. $\Box$
Proof of Theorem 18. Pictures ($a$) and ($b$) of Fig. 3 are realized, respectively, by the polynomials (22) and (23) with arbitrary $t>0$.
(22) $x^{4}-6x^{2}y^{2}+y^{4}+t$
(23) $x^{4}-6x^{2}y^{2}+y^{4}-2t(x^{2}+y^{2})+\frac{1}{2}t^{2}$
These polynomials are invariant under the entire group $G_{0}$. Therefore, their virtual components are associated with single real connected components of the discriminant complement.
Pictures ($c$) and ($i$) are realized respectively by polynomials (24) and (25) with arbitrary $t>0$.
(24) $(x^{2}-3t)^{2}-6(x^{2}-3t)(y^{2}-t/2)+(y^{2}-t/2)^{2}+t^{4}$
(25) $x^{4}-6x^{2}(y^{2}+t)+(y^{2}+t)^{2}-t^{2}/2$
These polynomials are invariant under the reflections $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y),$ which generate a subgroup of index two in $G_{0}$. Thus, in each of these two cases there are no more than two connected components of the discriminant complement associated with the corresponding virtual component.
Pictures ($d$), ($e$), ($f$), ($g$), and ($h$) are realized by the equations (26), (27), (28), (29), and (30) for sufficiently small $\varepsilon>0$ and arbitrary $t>0$.
(26) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}+\varepsilon t^{3}x\right)$
(27) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}-\varepsilon t^{3}x\right)$
(28) ${\bf T}\left(x^{4}-6x^{2}y^{2}+t(x^{3}-\varepsilon^{4}t^{2}x)+(y^{2}-\varepsilon^{3}t^{2})^{2}\right)$
(29) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}-6\varepsilon txy^{2}-2\varepsilon^{4}t^{3}x+8\varepsilon^{3}t^{2}y^{2}\right)$
(30) ${\bf T}\left((x+0.7t)(x+1.3t)(x+6t)(x+7t)-6x^{2}y^{2}+y^{4}\right)$
These five polynomials are all constructed as follows. First we take the polynomial
$x^{4}-6x^{2}y^{2}+y^{4}+x^{3},$
whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at its center. Then, we perturb this singular point by adding very small terms of lower degrees. These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [24]. All the resulting polynomials (2630) are invariant under the reflection in the axis $y=0$. Thus, in each of these five cases, the virtual component is associated with no more than four real connected components of the discriminant complement. These four components differ for the topological reasons.
For the remaining nine virtual components, the convenient realizations of the corresponding real connected components can be obtained from the polynomials (22)–(30) via the operation (17). $\Box$
3. Components for $J_{10}$ singularities
3.1.
The canonical versal deformations of the singularities of the classes $J_{10}^{1}$ and $J_{10}^{3}$ consist, respectively, of polynomials
(31) $(x-Ay^{2})(x^{2}+y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}+3y^{4})y,\qquad A\in(-\infty,+\infty),$
and
(32) $(x-Ay^{2})(x^{2}-y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}-3y^{4})y,\qquad A\in(-1,1),$
with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, cf. [10].
Definition 7.
A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (31) or (32) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal quasihomogeneous part of $f$ (of degree six with weights $\deg x=2,\deg y=1$) consists of three different complex parabolas or two parabolas and one line in ${\mathbb{C}}^{2}$.
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively. The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consists of 10 elements.
Proposition 25 (see [23]).
If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials either are the same or are mapped to each other by the involution
(33) $f(x,y)\leftrightarrow f(x,-y)\ .$
In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components. $\Box$
In the following pictures, the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
3.2. Components for $J_{10}^{1}$ singularities
++-$(a)$
++-$(b)$
++-$(c)$
++-$(d)$
++-$(e)$
++-$(f)$
++-($\tilde{b}$)
++-($\tilde{c}$)
++-($\tilde{d}$)
++-($\tilde{e}$)
Figure 4. Non-discriminant perturbations of $J_{10}^{1}$ singularities
Proposition 26 (see §4.2 in [24]).
There are exactly ten topological types of zero-level sets of non-discriminant perturbations of $J_{10}^{1}$ singularities, see Fig. 4. $\Box$
Proposition 27 (see [18]).
There are exactly ten virtual components of the formal graph of type $J_{10}^{1}$. These virtual components are represented by polynomials, the topological shapes of whose zero-level sets are shown on Fig. 4. $\Box$
Theorem 28.
There are exactly thirteen connected components of the discriminant complement in the space of all polynomials of class $\Omega_{1}$. Each of the virtual components shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig. 4 is associated with a single real connected component. Each of the virtual components/topological types represented in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((33)).
Proof. The topological types shown in pictures ($a$), ($b$), ($c$), and ($d$) are represented respectively by the following polynomials:
$x(x^{2}+y^{4}+t),$
$(x-2t)((x+t)^{2}+y^{4}-t^{2}),$
$(x-2t)((x+t)^{2}+y^{4}-3ty^{2}+2t^{2}),$
$(x-y^{2}+2t)(x^{2}-2tx+y^{4}-5t^{2})-t^{3}/10$
with arbitrary $t>0$. These polynomials are all invariant under the involution (33). Therefore, the corresponding virtual components are associated with single connected components of the discriminant complement.
The realizations of pictures (b~)($\tilde${b}), (c~)($\tilde${c}), and (d~)($\tilde${d}) can be obtained from these realizations of pictures $(b),$ $(c)$, and $(d)$ by the involution
(34) $f(x,y)\mapsto-f(-x,y).$
All of them are also invariant under the involution (33).
No polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(f)$ can belong to the same connected component as its image f(x,y)f(x,-y) under the involution (33). Indeed, the polynomials (31) are of degree three in the variable $x$, so any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times. Therefore, the ovals placed on the different sides of the infinite component of the zero-level set cannot pass one above the other. On the other hand, the involution (34) moves the component of picture (f) to the same component as (13).
An invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ realizing the picture $(e)$ from its mirror image f(x,y)f(x,-y) is as follows. Four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take three arbitrary points inside the rightmost three ovals. There is a single parabola (or line) of the form $x=ay^{2}+by+c$ that passes through these three points. The restriction of the polynomial $f$ to this parabola is a polynomial of degree at most six in the coordinate $y$. This restriction has six roots, so it is of degree exactly six. This parabola cannot intersect the fourth oval, hence it passes either above or below it with respect to the $x$ coordinate. The position of this parabola relative to the fourth oval is independent of the choice of three points inside the ovals, and remains unchanged along the connected components of the discriminant complement. However, these relative positions are different for any polynomial $f(x,y)$ with the topological shape of the zero-level set shown in picture $(e)$ and for its mirror image f(x,y)f(x,-y). The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
3.3. Components for $J_{10}^{3}$ singularities
---++++++$(a)$
---++++++$(b)$
---++++++$(c)$
---++++++$(d)$
---++++++$(e)$
---++++++$(f)$
---++++++$(g)$
---++++++$(h)$
---++++++$(i)$
---++++++$(j)$
---++++++$(k)$
---++++++$(l)$
---++++++$(m)$
Figure 5. Perturbations of $J_{10}^{3}$
Proposition 29 (see [18]).
There are exactly 23 virtual components of the formal graph of type $J_{10}^{3}$. The shapes of the zero-level sets of some polynomials $f(x,y)$ representing thirteen of these components are shown in Fig. 5. In all these cases except those shown in pictures $(h)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((34)), represent different virtual components. $\Box$
Remark 30.
Twelve topological types shown in these pictures are given in Fig. 29 of [24]. Pictures $(d)$ and $(e)$ of Fig. 5 are topologically equivalent, however they represent different virtual components.
We will denote by (a~)($\tilde${a}), (b~)($\tilde${b}),…the pictures obtained from the corresponding pictures $(a)$, (b)(b),… of Fig. 5 by the transformation (34), that is, by the reflection in a horizontal line and simultaneous change of all signs.
Theorem 31.
There are exactly 33 connected components of the complement of the discriminant variety of the versal deformation ((32)). Each virtual component represented by one of ten pictures
(35) (h),(h~),(i),(i~),(j),(j~),(k),(k~),(g),(l)(h),\ ($\tilde${h}),\ (i),\ ($\tilde${i}),\ (j),\ ($\tilde${j}),\ (k),\ ($\tilde${k}),\ (g),\ (l)
[…]
Proof. The difference of all these 33 connected components follows from the topological considerations and Bezout’s theorem, see [18]. By Proposition 25, it remains to be shown that, in all cases not covered by the list (35), the involution (33) does not change the real connected component. The following polynomials (36), (37), (38), (39), are all invariant under this involution, so their virtual components are represented by single connected components of the complements of the discriminants.
The polynomial
(36) $(x-t)^{3}-(x-t)y^{4}+3(x-t)^{2}+5(x-t)y^{2}-4(x-t)-Ct^{3}$
with arbitrary $t>0$ realizes the picture $(a)$ for $C<6-\frac{14}{3}\sqrt{\frac{7}{3}}$, the picture $(b)$ for $C\in\left(6-\frac{14}{3}\sqrt{\frac{7}{3}},-\frac{1}{2}\right)$, and the picture $(d)$ for $C\in\left(-\frac{1}{2},0\right)$.
Picture $(m)$ is represented by the polynomials
(37) $x(x-y^{2}-t)(x+y^{2}+t),\quad t>0.$
The polynomials
(38) $(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+3t(x-t)y^{2}+Ct^{3}$
realize picture $(c)$ for any $C\in(0,1/2)$.
The polynomials
(39) $(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+5t(x-t)y^{2}-4t^{2}(x-t)+\frac{t^{2}}{10}y^{2}-0.35t^{3}$
realize picture $(f)$.
Figure 6.
To realize the picture $(e)$, we start with the polynomial
(40) $x(x+(y-2)^{2}-2)(x+(y+2)^{2}-2),$
whose zero-level set is shown in Fig. 6. Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal quasihomogeneous part will be of class $J_{10}^{3}$. The obtained polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values. It will also have two critical points in the complex domain. Then we add a very small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain. The zero-level set of the resulting polynomial will have the desired shape shown in picture (e) of Fig. 5. Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal quasihomogeneous part of the resulting polynomial will have the canonical form, see (32). Finally, we apply a parallel translation of the $x$ coordinate to kill the monomial $x^{2}$. The resulting polynomial realizes the picture $(e)$, and is invariant under the involution (33), in particular it has the form (32) with $\lambda_{2}=\lambda_{4}=\lambda_{7}=\lambda_{9}=0$. $\Box$
4. The cases $P_{8}$
4.1.
Two classes of function singularities $P_{8}^{1}$ and $P_{8}^{2}$ in three variables have the canonical normal form
(41) $x^{3}+y^{3}+z^{3}-3Axyz,\quad A\neq 1,$
with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$. The zero sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected components in ${\mathbb{R}P}^{2}$. Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to the plane with a single singular point at the origin. The zero set of a polynomial of class $P_{8}^{2}$ consists of such a surface and a surface that is homeomorphic to the standard quadratic cone.
It follows easily from Morse theory that the closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f_{\lambda}$ of a singularity (41) with $\mbox{\rm Ind}(f_{\lambda})$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, three handles, or the disjoint union of the projective plane and the sphere)).
The standard versal deformations of $P_{8}^{1}$ and $P_{8}^{2}$ singularities consist of the polynomials
(42) $f_{\lambda,A}=f_{A}+\lambda_{1}+\lambda_{2}x+\lambda_{3}y+\lambda_{4}z+\lambda_{5}xy+\lambda_{6}xz+\lambda_{7}yz,$
where $f_{A}$ is the corresponding normal form (41).
Definition 8.
A polynomial $f:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all different and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
Let $\Xi_{1}$ and $\Xi_{2}$ denote the parameter spaces ${\mathbb{R}}^{7}\times(-\infty,1)$ and ${\mathbb{R}}^{7}\times(1,\infty)$ of the deformations (42) of the functions (41) of classes $P_{8}^{1}$ and $P_{8}^{2}$.
4.2. The homological invariant of real and virtual components
Let $f_{0}$ be a function singularity of class $P_{8}^{1}$ or $P_{8}^{2}$ having the form (41). Let $D$ be a small neighborhood of the point $\{f_{0}\}$ in the space $\Xi_{1}$ or $\Xi_{2}$. Let $B$ be a sufficiently large (compared to $D$) closed ball in ${\mathbb{R}}^{3}$, and let $f_{\lambda}$, $\lambda\in D$, be a small perturbation of $f_{0}$. Define $T_{-}(\lambda)$ as a real number slightly lower than the minimum of 0 and the lowest real critical value of $f_{\lambda}$. Similarly, define $T_{+}$ as a number slightly higher than the maximum of 0 and the highest real critical value of $f_{\lambda}$. For any $\lambda\in D$, define three sets: $W_{\lambda,0}=f_{\lambda}^{-1}((-\infty,0])\cap B$, $W_{\lambda,-}=f_{\lambda}^{-1}((-\infty,T_{-}])\cap B,$ and $W_{\lambda,+}=f_{\lambda}^{-1}((-\infty,T_{+}])\cap B$. The sets $W_{\lambda,-}$ and $B\setminus W_{\lambda,+}$ form then a trivializable fiber bundle over $D$. These sets contain their intersections with $\partial B$ as deformation retracts. In the case $P_{8}^{1}$ these intersections are just the discs. In the case $P_{8}^{2}$ each of them homeomorphic to the disjoint union of a disc and an annulus. In particular, the reduced homology groups $\tilde{H}_{i}(W_{\lambda,-})$ and $\tilde{H}_{i}(B\setminus W_{\lambda,+})$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
Definition 9.
The homology index $\mbox{HI}(f_{\lambda})$ of a non-discriminant function $f_{\lambda}$, $\lambda\in D$ is the triple of Betti numbers $b_{0},b_{1},b_{2}$ of the relative homology group $H_{*}(W_{\lambda,0},W_{\lambda,-})$.
It is convenient to consider this homology group in the context of the exact homological sequence of the triple
(43) $(W_{\lambda,+},W_{\lambda,0},W_{\lambda,-}).$
[…]
Clearly, the invariant Ind is equal to the alternating sum of three elements of the homology index.
4.3. The symmetry group and the relation between the real and virtual components
The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following exceptions: the frames of vanishing cycles in the 2-homology groups of zero-level sets of polynomials consist of eight elements, and the data set of virtual functions only includes the parities of the Morse indices of critical points, and not their integer values. We also have the following analog of Proposition 2
Proposition 32 (see [22], Corollary 8).
1. Any virtual Morse function of the type $\Xi_{1}$ or $\Xi_{2}$ is associated with a real polynomial of the same type.
2. If two virtual Morse functions of type $P_{8}^{1}$ (respectively, $P_{8}^{2}$) belong to the same virtual component, then they are associated with two generic Morse polynomials of class $\Xi_{1}$ (respectively, $\Xi_{2}$) that are isotopic to each other within the set of Morse polynomials. $\Box$
The group $S(3)$ of coordinate permutations in ${\mathbb{R}}^{3}$ preserves the normal form (41) and acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. This action is generated by the following transformations:
(44) $f(x,y,z)\mapsto f(y,x,z)\quad\mbox{and}\quad f(x,y,z)\mapsto f(y,z,x).$
The first transformation is an involution, and the second is of order three.
In particular, this group can act on the set of the connected components of the discriminant complement in the spaces $\Xi_{i}$. The transformations (44) preserve all topological characteristics of the functions. Thus, all connected components in the same orbit of this action are associated with the same virtual component of the formal graph of type $P_{8}^{1}$ or $P_{8}^{2}$.
Additionally, the involution
(45) $f(x,y,z)\mapsto-f(-x,-y,-z)$
also acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. It preserves the invariant Ind and defines involutions on both sets of real and virtual components of these classes.
Proposition 33.
If two generic polynomials, $f$ and $\tilde{f}$, of class $\Xi_{1}$ or $\Xi_{2}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of this class that contain $f$ and $\tilde{f}$ are mapped to each other by an element of the group $S(3)$.
Definition 10.
For any polynomial $f:{\mathbb{R}}^{3}\to{\mathbb{R}}$ of degree three with principal homogeneous part of the form (41), ${\bf T}(f)$ is the unique polynomial of the form (42) with the same principal homogeneous part, which is obtained from $f$ by a parallel translation in ${\mathbb{R}}^{3}$. Namely, it is the translation by the vector $\frac{1}{3}(a,b,c)$, where $a,b,$ and $c$ are the coefficients of $f$ at the monomials $x^{2}$, $y^{2}$, and $z^{2}$.
Lemma 34.
[…]
(46) ${\bf T}\left(x^{3}+y^{3}+z^{3}+6xyz+(x+y-2z)^{2}\right)$
and two other polynomials obtained from it by permuting the coordinates.
Proof. According to Lemma 35 of [22], there are only six polynomials of form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$. Three of these polynomials (mentioned in formula (26) of [22]) are locally equivalent to $\xi^{3}-\eta^{2}-\zeta^{2}+\frac{64}{3}$ at the $A_{2}$ point. The remaining three polynomials have the form (46), in particular belong to the same orbit of the group $S(3)$. $\Box$
Lemma 35.
The set of all polynomials of class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
(47) ${\bf T}\left(x^{3}+y^{3}+z^{3}-3(1+\sqrt{3})xyz+(x+y+(1+\sqrt{3})z)^{2}\right),\$
[…]
Proof. This follows from statement 2 of Lemma 44 of [22] (the coefficients $\varkappa$ and $c$ in formula (45) of [22] are fixed by our conditions on the critical values). $\Box$
Proof of Proposition 33. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 34 (respectively, 35) repeats the deduction of Proposition 8 from Lemma 9. $\Box$
4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
Proposition 36 (see [18], Proposition 7).
1. There are exactly 6503 virtual functions of the class $P_{8}^{1}$.
2. There are exactly seven virtual components of the formal graph of type $P_{8}^{1}$: two with $\mbox{\rm Ind}=1$, two with $\mbox{\rm Ind}=-3$, and one virtual component with each value 0, 1-1, and 2-2 of the Ind invariant.
[…]
Theorem 37.
Each virtual component of type $P_{8}^{1}$ is associated with only one connected component of the space of non-discriminant polynomials of class $\Xi_{1}$. So, there are exactly seven such connected components.
Proof. The invariant Ind of the polynomials
(48) $x^{3}+y^{3}+z^{3}-3(x+y+z)+C$
is equal to 0 for C<6C<-6 and $C>6$, to $1$ for $C\in(-6,-2)$ and $C\in(2,6)$, and to 2-2 for $C\in(-2,2)$. These polynomials are invariant under the action of the group $S(3)$ of permutations of coordinates. Therefore, all virtual components represented by these polynomials are associated with single real connected components of the set of non-discriminant polynomials.
Two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2. For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$. According to the previous paragraph, the associated virtual components are also different.
The polynomials
(49) ${\bf T}\left(x^{3}+y^{3}+z^{3}+3(x+y+z)^{2}-(x^{2}+y^{2}+z^{2})/2+\frac{7}{36}\right)$
and
(50) ${\bf T}\left(x^{3}+y^{3}+z^{3}-3(x+y+z)^{2}+(x^{2}+y^{2}+z^{2})/2-\frac{7}{36}\right)$
have $\mbox{Ind}=-3$. They are related to each other by the involution (45). Hence, by the last statement of Proposition 36, they realize two different virtual components. Since these polynomials are invariant under the entire group $S(3)$, these virtual components are also associated with single real components.
[…]
Lemma 38.
For each linear function $l\in L$,
1)) the polynomial
(51) $x^{3}+y^{3}+z^{3}-l^{2}$
has only one real critical point, namely a critical point of class $D_{4}^{+}$ at the origin. In particular, this polynomial has four critical points ((counted with multiplicities)) in the complex domain.
2)) For any sufficiently small $\varepsilon>0$, the polynomial
(52) $F(l)\equiv x^{3}+y^{3}+z^{3}-l^{2}-\varepsilon p_{l}^{2}$
has only two critical points in ${\mathbb{R}}^{3}$: a critical point of type $A_{3}$ at the origin, at which this polynomial is equal to
(53) $\xi^{4}-\eta^{2}-\zeta^{2}$
in some local coordinates, and a Morse critical point with Morse index $1$ and negative critical value.
Proof. 1. At each critical point of the polynomial (51), the sum of its three partial derivatives is equal to $3(x^{2}+y^{2}+z^{2}),$ so the origin is the only critical point of (51) in the real domain. In coordinates $l,p_{l},q_{l}$, this polynomial has the form
(54) $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}+l\psi_{2}(p_{l},q_{l})+l^{2}\psi_{1}(p_{l},q_{l},l),$
where $\psi_{2}$ and $\psi_{1}$ are some homogeneous polynomials of degrees 2 and 1, respectively.
The lowest quasihomogeneous part of the polynomial (54) in the coordinates $l,p_{l},$ and $q_{l}$ with weights $\deg l=3$ and $\deg p_{l}=\deg q_{l}=2$ is equal to $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}.$ Therefore, by the techniques of [2], §12, this polynomial has the normal form $Z^{2}+X^{3}+XY^{2}$ of $D_{4}^{+}$ class in some local coordinates.
2. Let $\alpha$ be the coefficient of the polynomial $\psi_{2}$ at the monomial $q_{l}^{2}$. Then the lowest quasihomogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$. The substitution $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2},P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial to the polynomial with the lowest quasihomogeneous part
$\tilde{l}^{2}-\varepsilon P_{l}^{2}+\left(\frac{\alpha^{2}}{4}+\frac{1}{4\varepsilon}\right)q_{l}^{4}\ .$
Again, using the techniques of §12 of [2], this function can be reduced to the standard form (53) of $A_{3}$ singularities via a local change of coordinates. If $\varepsilon$ is sufficiently small, then adding $\varepsilon p_{l}^{2}$ to (51) does not bring four non-real critical points of (51) back to the real domain. On the other hand, by the index considerations the obtained polynomial $F(l)$ has one Morse critical point in addition to the $A_{3}$ point at the origin. This Morse critical point depends continuously on $l$, therefore its Morse index is the same for all $l$. It is easy to calculate that for $l\equiv x$ it is equal to 1. Our critical point of class $A_{3}$ is topologically equivalent to a Morse critical point with Morse index 2 (and can be perturbed to such a point by an arbitrarily small change of the polynomial $F(l)$). The Morse-like complex of the polynomial $F(l)$ is acyclic, therefore the critical value of this $A_{3}$ type critical point should be higher than that of the only other critical point of Morse index 1. $\Box$
Denote by $C(l)$ the critical value of the polynomial $F(l)$ at its Morse critical point. Then the polynomial
$\tilde{F}(l)\equiv F(l)-\frac{C(l)}{2}$
is non-discriminant and has $\mbox{Ind}=-1$.
The polynomial $\tilde{F}(l)$ is canonically determined by the linear form $l$ and depends continuously on it. The set of all polynomials $\tilde{F}(l),$ $l\in L$, is connected and invariant under the action of the group $S(3)$. Thus the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement. $\Box$
4.5. One-dimensional cohomology of discriminant complements of $P_{8}^{1}$ singularities
Theorem 39.
The first cohomology group of the component of the discriminant complement of the $P_{8}^{1}$ singularities that consists of polynomials with $\mbox{Ind}=-1$ is non-trivial.
Proof. The zero-level set of the polynomials in this component is homeomorphic to the plane with a handle. Indeed, the topology of these sets remains constant along the component. For any polynomial $\tilde{F}(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to the plane) by passing through a Morse critical point of Morse index 1 once.
In particular, the first homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$. These homology groups form a local system over the component, in particular the fundamental group of the component acts on them by monodromy. For each $l\in L$, one basic homology class of the zero-level set of $\tilde{F}_{l}$ is defined by the intersection of this set and the plane $\ker l$. When $l\in L$ runs a half of the circle $L$, the plane $\ker l$ runs a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomial $\tilde{F}(l)$ runs a closed cycle in the discriminant complement. The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation. $\Box$
4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularity
Definition 11.
The invariant Card of a virtual component is the number of its vertices.
The involution (45) preserves this invariant.
Table 3. Statistics for $P_{8}^{2}$ singularity
No. 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
Ind 3-3 2-2 1-1 0 1
Card 258 156 60 60 1216 336 336 1318 844 844 1648 1648 262 94 94
Proposition 40 (see [18], Proposition 8).
1. There are exactly 9174 distinct virtual functions of $P_{8}^{2}$ singularity class.
2. These virtual functions are distributed over exactly fifteen virtual components. The values of invariants Ind and Card of these components are shown in the second and third rows of Table 3.
3. All pairs of different virtual components of $P_{8}^{2}$ singularity with the same invariant Card value can be realized by generic polynomials connected by involution ((45)).
Theorem 41.
Each of the ten virtual components of the formal graph of type $P_{8}^{2}$ that are mentioned in columns 1, 3, 4, 6, 7, 11, 12, 13, 14, and 15 of Table 3 is associated with only one real connected component of the discriminant complement.
Proof of this theorem takes the rest of this section.
All of the polynomials given in the next formulas (55), (56), (57), and (58) are invariant under the $S(3)$ action. Thus, for any polynomial $f$ of this list and any noncritical value $C$ of $f$, the set-theoretical invariant of the polynomial fCf-C is a virtual component associated with only one real component of the discriminant complement. We will show that these virtual components, or those obtained from them via the involution (45), include all virtual components with Card equal to 258, 336, 1648, 262, and 94. The proof of Theorem 41 for the components with $\mbox{Card}=60$ is more implicit, see § 43.
Components with $\mbox{\rm Ind}=0$ and $\mbox{\rm Card}=1648$. For each polynomial $f$ of class $\Xi_{2}$, the polynomials $f+C$ and fCf-C with sufficiently large $C$ realize two different virtual components with $\mbox{Ind}=0$. Indeed, the homology index of one of these components is $(000)$, and the other’s is $(011)$. In particular, the $S(3)$-invariant polynomials
(55) $x^{3}+y^{3}+z^{3}-4xyz+C\qquad\mbox{and}\qquad x^{3}+y^{3}+z^{3}-4xyz-C$
represent these two components for any $C>0$. These polynomials are not Morse, but any their sufficiently small generic perturbations are associated with virtual functions with $\mbox{Card}=1648$.
Components with $\mbox{\rm Card}=262$. The polynomial
(56) $x^{3}+y^{3}+z^{3}-4xyz+(x+y+z)$
has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2. Thus, it has invariant values $\mbox{Ind}=1$ and $\mbox{HI}=(00\bar{1})$. This polynomial is invariant under the involution (45). According to the last statement of Proposition 40, it cannot represent a component of the discriminant complement with $\mbox{Card}=94$. Therefore, it belongs to the component with $\mbox{Card}=262$.
Components with $\mbox{\rm Card}=94$ and $336$. The polynomial
(57) $f\equiv x^{3}+y^{3}+z^{3}-6xyz-\frac{3}{2}(x^{2}+y^{2}+z^{2})$
has eight real Morse critical points: one local maximum with critical value 0, three points of Morse index $2$ with value $\frac{11}{49}$, three points of index $1$ with value $\frac{1}{2}$, and one more point of index $2$ with value $\frac{3}{2}$. Therefore, the index $\mbox{Ind}(f-C)$ is equal to 0 for $C<-\frac{3}{2}$, to 1 for $C\in\left(-\frac{3}{2},-\frac{1}{2}\right),$ to 2-2 for $C\in\left(-\frac{1}{2},-\frac{11}{49}\right),$ to $1$ for $C\in\left(-\frac{11}{49},0\right),$ and to 0 for $C>0$.
The group $H_{*}(W_{\lambda,+},W_{\lambda,0})$ corresponding to the polynomials fCf-C with $C\in\left(-\frac{11}{49},0\right)$ is isomorphic to $\mathbb{Z}$} in dimension three and is trivial in all other dimensions. Therefore, by the exact sequence of the triple (43), the homology index $\mbox{HI}(f+\frac{1}{7})$ is equal to $(0,1,2)$. The involution (45) moves this polynomial to that with $\mbox{HI}=(1,0,0)$. In particular, these two polynomials represent two different components, and the associated virtual components are also different. According to Table 3, these are the components with $\mbox{Card}=94$.
The homology index $\mbox{HI}(f-C)$ with $C\in\left(-\frac{3}{2},-\frac{1}{2}\right)$ is $(0,0,1),$ because the corresponding group $H_{*}(W_{\lambda,0},W_{\lambda,-})$ is generated by the class of one critical point of Morse index $2$. Therefore, it is again the component with $\mbox{Card}=262$.
[…]
Components with $\mbox{\rm Card}=258,$ $336,$ and $94$. The polynomial
(58) $f=x^{3}+y^{3}+z^{3}-12xyz-\frac{1}{2}(x^{2}+y^{2}+z^{2})$
has eight real Morse critical points: a local maximum with critical value 0, three points with Morse index 2 and critical value $\frac{11}{3969}$, one point more with Morse index 2 and critical value $\frac{1}{162}$, and three points with index 1 and critical value $\frac{1}{54}$. Thus, $\mbox{Ind}(f-C)$ is equal to 0 for $C<-\frac{1}{54}$, to 3-3 for $C\in\left(-\frac{1}{54},-\frac{1}{162}\right)$, to 2-2 for $C\in\left(-\frac{1}{162},-\frac{11}{3969}\right)$, to $1$ for $C\in\left(-\frac{11}{3969},0\right)$, and to 0 for $C>0$.
Lemma 42.
[…]
Proof. By Proposition 40 of [22], the intersection matrix of vanishing cycles of the canonical basis for all generic polynomials of class $\Xi_{2}$ with eight real critical points, one of which is an extremum, is expressed by the Dynkin diagram shown in
Figure 7.
Fig. 7. This diagram does not contain edges of multiplicity 2-2. Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles is equal to 2-2. Therefore, the virtual component associated with our polynomial is the unique remaining virtual component with $\mbox{Ind}=-3$. $\Box$
Remark 43.
It is easy to see that the polynomials (58) with $\mbox{Ind}=-2$ and $\mbox{Ind}=1$ give new realizations of components with $\mbox{Card}=336$ and $94$, respectively.
Components with $\mbox{Card}=60$.
[…]
Theorem 44.
There exists a $S(3)$-invariant polynomial $f\in\Xi_{2}$ realizing a virtual component with $\mbox{\rm Card}=60.$
Proof. One of the two virtual components, each of which has $\mbox{Card}=60$, contains a virtual function that models a generic polynomial $f_{0}$ of class $\Xi_{2}$ such that
1) all of its critical points are real,
2) the three highest critical values are positive and the other five values are negative,
3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is shown in (59),
4) the Morse indices of the points 1, 2, 3, and 5 are odd, while the Morse indices of points 4, 6, 7, and 8 are even.
According to Proposition 32, there is a real polynomial $f_{0}\in\Xi_{2}$ with these data. According to Proposition 11 of [22], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), all odd Morse indices are equal to 1, and all even Morse indices are equal to 2.
Let $i:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $i(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by their increase, and $i(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$. Using the Lyashko–Looijenga map (in the form of [10]) we can lift the path $i$ uniquely into a path $I:[0,1]\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $i(t)$ is the sequence of critical values of the polynomial $I(t)$. As in the proof of Theorem 7 of [22], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined. Indeed, according to Proposition 2 of [10], this is guaranteed by the fact that all quadratic forms corresponding to the limit critical values are elliptic. The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with value $c_{-}$ and three critical points with value $c_{+}$.
Lemma 45.
This polynomial $I(1)$ is $S(3)$-invariant.
Proof. The critical points of the polynomial $I(1)$ are ordered according to the previous ordering of the critical points of the polynomial $f_{0}$, from which they were obtained via continuous deformation along the path $I$. This ordering can be uniquely extended to a continuous system of orderings of the critical points of each polynomial $\tilde{f}$ from a sufficiently small neighborhood $U$ of the point $I(1)$ in $\Xi_{2}$. Define the map $\Lambda:U\to{\mathbb{R}}^{8}$, mapping each point $\tilde{f}\in U$ to the sequence of its correspondingly ordered critical values. By the versality property, this map is a local diffeomorphism. Let $\varepsilon>0$ be a very small number. For any permutation $\sigma\equiv\binom{1\ 2\ 3}{i_{1}\,i_{2}\,i_{3}}$ of critical points of $I(1)$ with critical value $c_{-}$, there is a unique point $\tilde{f}_{\sigma}\in U$, whose critical values at the critical points with numbers $i_{1},4,5,$ and $5+i_{1}$ are the same as those of $I(1)$, the values at the points $i_{2}$ and $5+i_{2}$ are equal to $c_{-}+\varepsilon$ and $c_{+}+\varepsilon$, and the values at the points $i_{3}$ and $5+i_{3}$ are equal to $c_{-}-\varepsilon$ and $c_{+}-\varepsilon$, respectively. Different permutations $\sigma$ of critical points define six different points $\tilde{f}_{\sigma}\in U$. All of these points are associated with the same virtual function, because any permutation $\sigma$ of the first three rows and columns in (59) and the simultaneous permutation of the last three rows and columns by the same rule $\sigma$ preserves the matrix.
As in the proof of Proposition 33 (repeating the proof of Proposition 8), the six points $\tilde{f}_{\sigma}\in U$ are all mapped to each other by the action of elements of the group $S(3)$. Therefore, each such element maps a point from an arbitrarily small neighborhood $U$ of the polynomial $I(1)$ to another point from the same neighborhood. By continuity, it also maps the polynomial $I(1)$ to itself. $\Box$
By the construction, the Card invariant of the virtual function associated with arbitrarily small perturbations of the polynomial $I(1)$ is equal to 60. This proves Theorem 44. $\Box$
The involution (45) moves the polynomial $I(1)$ to a polynomial in the other component of the discriminant complement with $\mbox{Card}=60$. Thus, this component is also represented by an $S(3)$-invariant polynomial. This concludes the proof of Theorem 41. $\Box$
5. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at parabolic singularities provides a complete list of local Petrovskii lacunas at the corresponding singular points of wavefronts of hyperbolic PDEs. In this theory, generating families of functions play a crucial role. Wave functions are defined on the parameter spaces of these families via integrals along certain cycles in the zero-level sets of the corresponding functions. Wavefronts are the discriminant sets of these families. Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, and the wave function is regular as well. However, over the discriminant sets of parameters, the level sets containing the integration cycles degenerate and the wave function may be singular. The singularity of a wave function in a neighborhood of a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point. The type of this asymptotic behavior depends primarily on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle. The homology classes depend on the wavefront complement components from which we approach the wavefront point. Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
If the singularity of the discriminant function corresponding to a wavefront point is isolated, then whether a neighboring component of the discriminant complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary function from that component (see, for example, [16]).
Versal deformations of stably equivalent function singularities can be naturally identified with each other. This identification establishes a one-to-one correspondence between the components of their discriminant complements. The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number of variables, and on the parities of the inertia indices of the quadratic part of the singularity. For example, the functions $f(x_{1},\dots,x_{n})$ and $\tilde{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}+x_{n+2}^{2}$ have the same set of local lacunas. On contrary, the function $\check{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}-x_{n+2}^{2}$ or the stably equivalent functions in n+2k1n+2k-1 variables generally has a different set.
Theorem 46.
The numbers of local lacunas near parabolic singularities of wavefronts in versal generating families are given in Table 4.
Table 4. Numbers of local lacunas at parabolic singularities
Singularity NN even NN even NN odd NN odd
class $i_{+}$ even $i_{+}$ odd $i_{+}$ even $i_{+}$ odd
$P_{8}^{1}$ 0 0 22 0
$P_{8}^{2}$ 0 0 33 0
$X_{9}^{+}$ 11 0 22 0
$X_{9}^{-}$ 11 0 0 22
$X_{9}^{1}$ 0 0 0 0
$X_{9}^{2}$ 0 44 0 0
$J_{10}^{3}$ 0 11 0 0
$J_{10}^{1}$ 0 0 0 0
In comparison with the results of [17], this table replaces the lower estimate $\geq 2$ for $P_{8}^{2}$ with the exact number 3, and replaces some five estimates of the form $\geq k$ with the same numbers $k$.
In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity for even $N$ and $i_{+}$ is component $(a)$ of Fig. 1. The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure. The four lacunas at the $X_{9}^{2}$ singularity with even $N$ and odd $i_{+}$ are shown in picture $(i)$ of Fig. 3, or in its rotation by the angle $\frac{\pi}{2}$, or in the pictures obtained by applying the operation (17) to these two. The unique lacuna at the $J_{10}^{3}$ singularity with even $N$ and odd $i_{+}$ is shown in picture $(m)$ of Fig. 5. The two local lacunas of the $P_{8}^{1}$ singularity at odd $N$ and even $i_{+}$ are the components with $\mbox{Ind}=1$. Three local lacunas of the $P_{8}^{2}$ singularity with odd $N$ and even $i_{+}$ are also the components with $\mbox{Ind}=1$. The only one new local lacuna found now is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.